### v2

# SweetTokenizer: Semantic-Aware Spatial-Temporal Tokenizer for Compact Visual Discretization

Abstract
This paper presents the Semantic-aWarE spatial-tEmporal Tokenizer (SweetTokenizer), a compact yet effective discretization approach for vision data. Our goal is to boost tokenizers’ compression ratio while maintaining reconstruction fidelity in the VQ-VAE paradigm. Firstly, to obtain compact latent representations, we decouple images or videos into spatial-temporal dimensions, translating visual information into learnable querying spatial and temporal tokens through a Cross-attention Query AutoEncoder (CQAE). Secondly, to complement visual information during compression, we quantize these tokens via a specialized codebook derived from off-the-shelf LLM embeddings to leverage the rich semantics from language modality. Finally, to enhance training stability and convergence, we also introduce a curriculum learning strategy, which proves critical for effective discrete visual representation learning. SweetTokenizer achieves comparable video reconstruction fidelity with only 25% of the tokens used in previous state-of-the-art video tokenizers, and boost video generation results by 32.9% w.r.t gFVD. When using the same token number, we significantly improves video and image reconstruction results by 57.1% w.r.t rFVD on UCF-101 and 37.2% w.r.t rFID on ImageNet-1K. Additionally, the compressed tokens are imbued with semantic information, enabling few-shot recognition capabilities powered by LLMs in downstream applications.
1 Introduction
Figure 1: Illustration of our framework. We build a compact visual latent space by reducing token count and leveraging semantic text embedding. The encoded tokens can be applied to downstream tasks, such as generation and understanding.
Visual tokenizers [41, 6, 45, 43, 12, 53, 55] are emerging as essential components in the field of modern computer vision models, particularly in the generation [12, 55, 50] and understanding [44, 18, 36, 28, 47] of vision data. These tools convert visual inputs into discrete tokens, capturing essential temporal and spatial features that facilitate advanced analysis by formulating visual-related tasks as a token prediction process.
Compression ratio and reconstruction fidelity are vital criteria for evaluating a tokenizer. However, recent visual tokenizers, especially video tokenizers [45, 43, 55] typically retain a low compression ratio. This is because visual tokens are usually derived from 2D patches [10] or 3D tubes [45, 12] which preserve location relationships (e.g., each token corresponds to a specific region of input [56]), leading to redundancy in both spatial and temporal dimensions. To reduce token count, we take inspiration from Q-Former [27], which compresses raw input into learnable “query tokens”. However, it is observed that directly flattening video tokens into sequence may lead to catastrophic temporal information loss. Therefore, we explore a spatial-temporal decoupled approach to perform tokenization.
A higher compression ratio generally makes reconstruction more challenging. To complement visual information during compression, we follow [29, 54, 57] to introduce pretrained language embeddings as the latent codebook, leveraging the powerful semantic representation capabilities of large language models (LLM) [1, 38]. However, previous works primarily focus on image modality, overlooking the relationships between text and motion in video domain.
To address existing limitations, we propose SweetTokenizer – Semantic-aWarE spatial-tEmporal Tokenizer – as illustrated in Figure 1. First, considering the heterogeneous redundancy in static images and dynamic frames, we propose the Cross-attention Query AutoEncoder (CQAE) to compress spatial and temporal information into separate learnable queries. Different from previous works [56, 27], our findings indicate that coupling the compression of spatiotemporal information increases the difficulty for the decoder to learn the motion information of the same pixel across consecutive frames. Thus, taking the decoupled spatial and temporal queries as inputs, we devise a strategy of spatial decoding followed by temporal decoding to achieve a separate reconstruction of the spatial and temporal dimensions of visual information. Additionally, the decoupled spatiotemporal reconstruction approach allows for pre-training with a large amount of image data, enhancing the model spatial representation. Second, to integrate the semantic information inherent in large language model (LLM), we design specialized codebooks tailored for spatial and temporal compression addressing the differences in semantic representation between spatial and temporal information. Specifically, we design two language-based codebooks based on the part of speech, using nouns and adjectives for spatial static information and verbs and adverbs for temporal motion information. By incorporating language-based codebooks, the learnable compressed queries can also be easily adapted to downstream visual understanding tasks by in-context learning of LLM. Third, we introduce the curriculum learning [2] mechanism by training SweetTokenizer in three stages to achieve stable convergence. We initially pre-train the spatial CQAE with image data, then joint training of the spatial and temporal CQAE with video data. The first two stages are supervised by the commitment loss of the codebook and the cross-entropy loss of the proxy code [56] from the pretrained tokenzier. In the last stage, we train the whole network and introduce the spatiotemporal decoder to reconstruct the original video.
Exhaustive experiments demonstrate the effectiveness of SweetTokenizer. Compared with SOTA methods [45], SweetTokenizer achieves comparable rFVD on the UCF-101 [34] while utilizing only 25% of the tokens. Notably, it enhances the gFVD metric from 191 to 128. At an equivalent compression rate on UCF-101, SweetTokenizer reduces rFVD from 42 to 18 and gFVD from 191 to 154. On the ImageNet-1k [8], SweetTokenizer demonstrates a substantial improvement in rFID, decreasing it from 0.59 to 0.37.
In summary, our work makes the following key contributions:
  • We introduce SweetTokenizer, a cutting-edge visual tokenizer that achieves decent reconstruction fidelity with a high compression ratio via spatial-temporal decoupling and cross-attention query autoencoder, reaching a “sweet spot” between compression and fidelity.
  • We leverage a semantic-enhanced latent codebook to utilize the off-the-shell representation capability of LLM embeddings, which improves reconstruction quality and facilitates downstream video understanding tasks.
  • We design a progressive training schedule following the curriculum learning mechanism, which ensures better convergence of visual tokenizers.
  • We perform extensive experiments to verify the effectiveness of SweetTokenizer, which exhibits the state-of-the-art performance on video reconstruction, image reconstruction, and class-conditional video generation tasks, leading by a large margin of 57.1%, 37.2%, and 32.9%.
Figure 2: Main pipeline. Our semantic-aware spatial-temporal tokenizer (SweetTokenizer) is trained in a vector-quantized variational autoencoder (VQ-VAE) style. We decouple video into spatial-temporal domains, then encode/decode video via the proposed novel cross-attention query autoencoder (CQAE), which significantly reduces the token number from 5,120 to 1,280 (Section 3.2.2). We further enhance our model with a language-based latent codebook, improving reconstruction fidelity and adding extra semantics (Section 3.2.3). A curriculum learning strategy is adopted to facilitate our training in three stages: Stage 1 and Stage 2 use loss marked purple; Stage 3 uses loss marked blue (Section 3.2.4). The quantized spatial-temporal latent queries $\hat{\Omega}_{s},\hat{\Omega}_{t}$ are used for downstream tasks.
2 Background
2.1 Visual Tokenizer With Vector Quantization
Exploring visual tokenizers and their applications in generative models has led to significant advancements in image/video-related tasks. The general idea is to discretize visual data into tokens, then tasks like visual generation [6, 51, 52, 12] & understanding [10, 4, 44, 19, 18, 36, 28, 47] can be tackled in a sequence prediction style as natural language processing [9, 30, 38]. Our work belongs to the series of Vector Quantized Variational AutoEncoder (VQ-VAE) [41, 32] tokenizers, which introduce a discrete latent space for continuous VAE [22] encoder-decoder structure. It typically encodes a high-dimensional image into a low-dimensional latent representation, then queries the nearest index from a learnable codebook to quantize the latent vector, and finally decodes back reversely to reconstruct the raw input signal. Since this type of tokenizer acquires reconstruction loss, it can maintain high-level semantic and low-level details of input vision. VQGAN [11] adopted adversarial training loss to improve high-frequency details. ViT-VQGAN [51] upgraded encoder-decoder with vision-transformer (ViT) architecture [10] and further boosted results. TiTok [56] replaced 2D image structure with 1D sequence latent representation, then used a self-attention transformer [42] to compress token number.
However, the above methods can only process image data. For video modality, TATS [12] used 3D-CNN to encode video patches and adopted sliding windows to deal with long-term relations. CViViT [43] used ViT [10] structure to encode spatial patches and then adopted a causal transformer to model temporal information. OmniTokenizer [45] and MAGVIT [53, 55] adopted similar transformer architecture and introduced image pre-training to improve video tokenizer. In this paper, we inherit the popular spatial-temporal decomposition design for video data. Moreover, we significantly increase the token compression ratio with a novel stacked cross-attention query autoencoder, which is inspired by Q-Former [27]’s token compression mechanism.
2.2 Language-based Latent Codebook
The codebooks learned by vanilla VQ-VAEs are not interpretable with lexical meanings. Therefore, many works attempt to utilize pretrained language models embedding codebooks to enhance semantics. LQAE [29] replaced the visual codebook with frozen word embeddings from BERT [9]. SPAE [54] quantized image latent space in a pyramid structure to preserve semantic information from low-level to high-level. It also used large language model (LLM) codebook [7] so that the encoded image token can be directly adapted to visual understanding tasks through in-context learning [3] ability of LLM. We follow this evaluation pipeline for few-shot classification in our paper. V2L-Tokenizer [59] utilized CLIP [31] pretrained encoder and injected a learnable projector to align visual-text latent space implicitly. VQCT [57] replaced the projector with graph convolution networks [24] to consider the relationship between vocabularies. Furthermore, De-Diffusion [48] directly encoded image into plain text as latent space interface and decodes back through a text-to-image (T2I) diffusion model [33]. However, none of these works dives deeply into the codebook design for video modality. Therefore, we propose splitting the codebook according to the video’s spatial-temporal attribute, which uses nouns & adjectives for spatial information, and verbs & adverbs for temporal information. This design helps us better align visual-motion-text semantics.
3 Method
3.1 Preliminary
A typical visual vector-quantization (VQ) model [12, 45, 53, 55] contains three parts: encoder $\mathcal{E}$, decoder $\mathcal{D}$ and latent quantizer $\mathcal{Q}$. Take video modality as an example, given a video input $x\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ represents the temporal length and $H\times W$ denotes spatial resolution, encoder $\mathcal{E}(x)$ projects it into latent space $\mathbb{Z}\in\mathbb{R}^{N\times D}$, where $D$ is latent dimension and $N$ is token number. A quantizer $\mathcal{Q}$ is constructed in this latent space $\mathbb{Z}$ by querying the nearest neighbor in codebook $C\in\mathbb{R}^{L\times D}$, where $L$ is codebook size. Then $\mathcal{D}$ decodes latent space back to pixel space and applies self-supervised reconstruction loss:
$\mathcal{L}_{rec}(x,\mathcal{D}(\mathcal{Q}(\mathcal{E}(x)))).$ (1)
We aim to learn a more compact latent space $\mathbb{Z}$ while maintaining the reconstruction fidelity. To this end, we design a semantic-enhanced spatial temporal decoupled tokenizer (SweetTokenizer) to address the issue of token compression (Figure 2). The main component is Cross-attenion Query AutoEncoders (CQAE), which aggregates information from disentangled spatial-temporal dimensions into learnable queries, achieving a 4$\times$ compression ratio improvement over previous video modality tokenizers. To enhance latent space representation and reconstruction fidelity, we incorporate compact text embeddings and refine vocabularies by appearance/motion attributes for better semantic alignment. Finally, we propose a progressive training schedule to stabilize convergence.
3.2 Spatial-Temporal Tokenization
3.2.1 Patchify
Given a video frame sequence $x\in\mathbb{R}^{T\times H\times W\times 3}$, we select the first frame $x_{0}$ as a reference for spatial information, the remaining $T-1$ frames $x_{1:T}$ for temporal information, following the strategy in [45]. We apply two patch kernels with shapes $p_{h}\times p_{w}$ and $p_{t}\times p_{h}\times p_{w}$ to $x_{0}$ and $x_{1:T}$ separately, generating $v_{s}\in\mathbb{R}^{1\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ and $v_{t}\in\mathbb{R}^{\frac{T-1}{p_{t}}\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ as inputs for transformer-based autoencoder, where $v_{s}$ contains spatial information, $v_{t}$ contains temporal information, with $D$ representing the hidden embedding size. In practice, for a video with 17 frames and a resolution of $256\times 256$, we set $(p_{t},p_{h},p_{w})$ to (4, 8, 8), thus patchify frames into $v_{s}$ with shape $1\times 32\times 32$ and $v_{t}$ with shape $4\times 32\times 32$. In Figure 2, we use $t=\frac{T-1}{p_{t}}=4$ to denote $v_{t}$’s length.
Figure 3: Module structure of Cross-attention Query AutoEncoder (CQAE). Information is transported from $K,V$ head to $Q$ head through stacked CQAE transformer layers. The enlarged view illustrates the detailed architecture.
3.2.2 Cross-attention Query AutoEncoder (CQAE)
So far, the raw video has been split into patches with a fixed ratio $p_{t}\times p_{h}\times p_{w}$. However, many works [56, 27, 37] have demonstrated that visual information remains redundant spatially and temporally, even after tokenization. To select “key words” from these tokens, we adopt a strategy similar to Q-Former [27], compressing patch sequence into a fixed number of query tokens via cross-attention interactions. As an innovation, we recursively inject these cross-attention query modules into transformer-based autoencoder to transfer information, forming our CQAE module. Moreover, we show in experiments that naively flattening patches into a long 1D sequence is unsuitable for video modality since it may lose temporal motion information in consecutive frames. Therefore, we introduce two autoencoders, $CQAE_{s}$ and $CQAE_{t}$, along with two sets of latent query token embeddings $\Omega_{s}\in\mathbb{R}^{N_{s}\times D}$ and $\Omega_{t}\in\mathbb{R}^{N_{t}\times D}$, for summarizing spatial and temporal information, where $N_{s}$ and $N_{t}$ control the latent token number. We set $N_{s}$ to 256 and $N_{t}$ to 1,024, resulting in 1,280 tokens in total.
During the encoding stage, we pass $v_{s}$ into $CQAE_{s}$ and $v_{t}$ into $CQAE_{t}$. Take $CQAE_{s}$ as example, as shown in Figure 3, $v_{s}$ is first reshaped from $1\times 32\times 32$ into $1\times 1024$, then passed through $K$-layer self-attention transformer to produce $K$ intermediate hidden units $[h(v_{s}^{1}),h(v_{s}^{2}),...,h(v_{s}^{K})]$. Meanwhile, the latent query token $\Omega_{s}$ is also passed through self-attention transformers, except that each hidden unit $h(\Omega_{s}^{i})$ is served as query tensor for cross-attention layer with $h(v_{s}^{i})$. Therefore, we gradually transfer spatial information from $v_{s}$ into $\Omega_{s}$ through interactions between hidden units $h(v_{s}^{i})$ and $h(\Omega_{s})$, reducing the spatial token number from 1,024 to 256.
Similarly, $v_{t}$ is processed with $CQAE_{t}$. Instead of directly using $v_{t}$, we employ frame-wise residual $\Delta v_{t}$ to generate temporal hidden units $[h(\Delta v_{t}^{1}),h(\Delta v_{t}^{2}),...,h(\Delta v_{t}^{K})]$, as [18] found that frame motion residuals can be more effectively represented. Notably, we reshape $\Delta v_{t}$ from $4\times 32\times 32$ into $1024\times 4$, and $\Omega_{t}$ from $1024$ into $1024\times 1$ to align the transformer inputs shape. Thus, the temporal information is embedded into $\Omega_{t}$ from $v_{t}$ using the same cross-attention querying approach along the temporal axis. $\Omega_{s}$ and $\Omega_{t}$ are quantized into $\hat{\Omega}_{s}$ and $\hat{\Omega}_{t}$, which will be explained in Section 3.2.3 later.
During the decoding stage, we use mask tokens $\Omega_{m}\in\mathbb{R}^{1\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ to receive information from $\hat{\Omega}_{s}$ and $\hat{\Omega}_{t}$, similar to the mask token prediction tasks in [9, 14, 37, 56]. As shown in Figure 3, yellow blocks denote that the tensor contains spatial visual information, while grey blocks indicate that the tensor is information-free. We randomly initialize these mask tokens $\Omega_{m}$ and use them as a “canvas” to reconstruct the raw inputs by receiving visual information from $\hat{\Omega}_{s}$. The decoding stage consists of two steps. Firstly, the mask token $\Omega_{m}$ starts with shape $32\times 32$, which is reshaped into $1\times 1024$ to interact spatially with $\hat{\Omega}_{s}$. Each decoder layer has the same structure as the encoder layer, except that the query is $h(\Omega_{m}^{i})$ and the key/value is $h(\hat{\Omega}_{s}^{i})$. Afterwards, $\Omega_{m}$ is expanded from $1\times 1024$ into $5\times 1024$, in order to interact temporally with $\hat{\Omega}_{t}$. The final output $\Omega_{m}$ is a mixture of yellow and green block as illustrated in Figure 2, indicating that it simultaneously contains spatial and temporal information.
3.2.3 Language-based Latent Codebook
Previous works [29, 54, 57] have shown that text representations can enhance image VQ-VAEs , as the text provides additional semantic information from pre-trained language models. However, previous works mainly focus on the relationship between static image appearance and text semantics. Our experiment in Table 5 shows this is insufficient for video data, as static and motion information are typically embedded in different subsets of vocabularies.
To address this, we construct two separate codebooks, the spatial quantization codebook and the temporal quantization codebook, according to vocabulary attributes, as illustrated in Figure 2. We first extract candidate vocabularies from video captions, which is obtained by Qwen-2.5B [1] instruct model. We filter out words that occur with low frequency. Afterward, we extract Qwen-2.5B [1] text embedding of these vocabularies to fill in the columns of our codebook $C\in\mathbb{R}^{L\times D}$ and split it into four subsets: nouns, adjectives, verbs, and adverbs. Note that $C_{noun},C_{adj}$ are for spatial latent tokens $\Omega_{s}$ while $C_{verb},C_{adv}$ are for temporal latent tokens $\Omega_{t}$.
Given two encoded continuous latent vectors: $z_{s}\in\Omega_{s}$ and $z_{t}\in\Omega_{t}$, $z_{s}$ is passed through spatial quantization codebook, and $z_{t}$ is passed through temporal quantization codebook. The quantized $\hat{z}_{s}$ and $\hat{z}_{t}$ are obtained by nearest neighbor searching:
$z_{s},z_{t}=\mathcal{E}(x),$ (2)
$\hat{z}_{s}=\mathcal{P}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{noun}\cup C_{adj}}||z_{s}-\mathcal{P}(c_{i})||,$ (3)
$\hat{z}_{t}=\mathcal{P}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{verb}\cup C_{adv}}||z_{t}-\mathcal{P}(c_{i})||.$ (4)
To maintain text semantic information, we freeze the codebook and use a tiny projector network $\mathcal{P}$ to map from text space into visual space, following the strategy in [57]. Finally, the gradient is passed to the encoder via vector-quantization commitment loss proposed in [41], a common method to approximate differentiability:
$\mathcal{L}_{vq}=||sg[z_{s}]-\mathcal{Q}(z_{s})||^{2}+||z_{s}-sg[\mathcal{Q}(z_{s})]||^{2}$ (5)
$+||sg[z_{t}]-\mathcal{Q}(z_{t})||^{2}+||z_{t}-sg[\mathcal{Q}(z_{t})]||^{2}$
where $sg[\cdot]$ is stop-gradient operator. Figure 4 also shows that the encoded latent words by SweetTokenizer capture semantic meanings related to both visual appearance and motion.
Figure 4: The semantics of spatial-temporal “words”. The attention weights of the last encoder’s cross-attention layer are visualized via heatmap, showing the visual regions corresponding to the related latent words.
3.2.4 Progressive Training Schedule
To ensure better convergence, we follow the philosophy of curriculum learning [2], thus designing a progressive training schedule comprising three stages.
We adopt the proxy codes strategy following [56] to leverage pre-trained visual priors. Assuming $\Psi_{e}(\cdot)$ is a pre-trained visual tokenizer and $\Psi_{d}(\cdot)$ is a pre-trained pixel decoder. $\Psi_{e}(\cdot)$ maps visual inputs $x$ into well-trained integer token indices $\Psi_{e}(x)$, which serves as a proxy code classification label of our CQAE model, similar to the knowledge distilling approach [16]. The motivation is that we could bypass the early oscillatory training phase of GAN loss [13] as demonstrated in [56, 11], which allows us to apply it later once the training becomes stable. In practice, we use OmniTokenizer [45]’s encoder and decoder as $\Psi_{e}(\cdot)$ and $\Psi_{d}(\cdot)$.
First, we train $CQAE_{s}$ on spatial patches $v_{s}$ using image data only as pre-training in Stage 1, with the supervision of proxy code:
$\mathcal{L}_{stage1}=\mathcal{L}_{c}(CQAE_{s}(v_{s}),\Psi_{e}(x))+\mathcal{L}_{vq},$ (6)
where $\mathcal{L}_{c}$ denotes cross entropy loss and $\mathcal{L}_{vq}$ denotes codebook commitment loss in Eq 5. Note that SweetTokenizer adopts spatial-temporal disentanglement, so the model weights of $CQAE_{s}$ can be later adapted to image-level tasks via finetuning. Experiments in Table 3 show that our spatial branch already achieves satisfactory image reconstruction results.
Second, we jointly train $CQAE_{s},CQAE_{t}$ on $v_{s},v_{t}$ using video data in Stage 2. Since $CQAE_{s}$ is well-initialized, $CQAE_{t}$ can focus on learning temporal motions:
$\mathcal{L}_{stage2}=\mathcal{L}_{c}(CQAE_{s}(v_{s}),\Psi_{e}(x))+$ (7)
$\quad\quad\quad\quad\mathcal{L}_{c}(CQAE_{t}(v_{t}),\Psi_{e}(x))+\mathcal{L}_{vq}.$
Finally, in Stage 3, we combine the pre-trained pixel decoder $\Psi_{d}(\cdot)$ together with our CQAE and finetune them by pixel-level losses in Stage 3 to boost final performance:
$\mathcal{L}_{stage3}=\mathcal{L}_{(m,g,p)}(\Psi_{d}(CQAE(v)),x),$ (8)
where $\mathcal{L}_{m}$ denotes mean-squared error loss, $\mathcal{L}_{g}$ denotes GAN loss [13] and $\mathcal{L}_{p}$ denotes perceptual loss [20].
4 Experiments
4.1 Experiments Settings
Tokenizer #Tokens rFVD $\downarrow$
UCF-101 K-600
MaskGIT [6] 4352 240 202
VQGAN [11] 4352 299 270
TATS [12] 4096 162 -
MAGVIT [53] 4096 58 -
OmniTokenizer [45] 5120 42 26
SweetTokenizer 5120 18 8
SweetTokenizer 256 + 1024 44 28
Table 1: Video reconstruction FVD on the UCF-101 and K-600 dataset, using a frame resolution 256 × 256. $\ast$” denotes training SweetTokenizer without token compression.
Tokenizer #Params #Tokens gFVD $\downarrow$
CogVideo [17] 9.4B 6800 626
TATS [12] 321M 4096 332
Video-LaVIT [18] 7B 512 280
OmniTokenizer [45] 650M 5120 191
SweetTokenizer 650M 5120 154
SweetTokenizer 650M 256 + 1024 128
Table 2: Comparison of class-conditional generation results on UCF-101. Each video is composed of 17 frames with a resolution of 256 $\times$ 256. We only compare autoregressive generators (AR) as baselines. “#Params” refers the parameter number of generator.
Dataset.
We evaluate the tokenization performance of SweetTokenizer on image and video datasets, including ImageNet [8], UCF-101 [34], and Kinetics-600 [21, 5]. Following [45], all images and video frames are resized to 256 × 256 resolution for experiments. The semantic capabilities of SweetTokenizer are tested through few-shot image classification on Real-Name Open-Ended miniImageNet [39] and few-shot video action recognition on UCF-101, as described in [58].
Evaluation Metrics.
For video reconstruction experiments, we evaluate using the Reconstruction Frechet Video Distance (rFVD) [40]. For video generation, we use the Generation Frechet Video Distance (gFVD) metric. For image reconstruction, we categorize recent methods by the number of compressed tokens, with each group assessed using the Frechet Inception Distance (FID) [15].
Tokenizer #Tokens Codebook Size rFID $\downarrow$
VQGAN [11] 256 1024 7.94
RQ-VAE [26] 256 16384 3.20
MaskGIT[53] 256 1024 2.28
LlamaGen-16 [35] 256 16384 2.19
TiTok [56] 256 4096 1.03
SweetTokenizer 256 10481 0.73
ViT-VQGAN [51] 1024 8192 1.28
OmniTokenizer [45] 1024 8192 1.11
OmniTokenizer [45] 1024 8192 0.69
LlamaGen-8 [35] 1024 16384 0.59
SweetTokenizer 1024 10481 0.37
Table 3: Image reconstruction FID on the ImageNet dataset, using a resolution of 256 $\times$ 256. $\diamond$” denotes continuous latent space without quantization. $\ast$” denotes training SweetTokenizer without token compression.
Implementation Details.
SweetTokenizer adopts a spatial-temporal architecture consisting of 8 spatial layers and 4 temporal layers, with both the encoder and decoder configured to a hidden dimension of 512. The latent space dimension is set to 256. For the LLM codebook quantizer, we exclude words with a frequency below 5, resulting in a selection of 5,078 nouns, 5,403 adjectives, 9,267 verbs, and 1,872 adverbs. This forms a spatial codebook of size 10,481 and a temporal codebook of size 11,139. We utilize a graph convolution network (GCN) to project Qwen-2.5 embeddings [1] into the visual latent space. Graph edges are constructed when a pair of words co-occur within a 5-token window in the text. The model is trained with a progressive three-stage schedule: Stage 1 uses a batch size of 32 for 300K iterations, Stage 2 uses a batch size of 8 for 300K iterations, and Stage 3 uses a batch size of 4 for 500K iterations. All training is performed on NVIDIA A100 GPUs. Adam [23] is employed for optimization ($\beta_{1}=0.9$ and $\beta_{2}=0.99$). During each stage, we use a cosine learning rate scheduler with a max learning rate of 1e-4 and a min learning rate of 1e-5, warmed up by 10K iterations.
OmniTokenizer [45]
SweetTokenizer (ours)
Figure 5: Class-conditional video generation result using OmniTokenizer [45] and our SweetTokenizer, on UCF-101 dataset. The action class label of each row is: “BlowingCandles”, “PlayingTabla”, “TennisSwing”, “PlayingViolin” and “HorseRiding”.
4.2 Video Reconstruction & Generation
We first evaluate the tokenization capability of SweetTokenizer on the UCF-101 and K-600 video datasets. As shown in Table 1, SweetTokenizer uses only 1,280 tokens (256 spatial tokens and 1,024 temporal tokens), which is four times fewer than OmniTokenizer’s 5,120 tokens. Despite this significant reduction, it achieves comparable performance, with rFVD scores of 44.35 on UCF-101 and 27.95 on K-600, demonstrating its effectiveness in compressing tokens while maintaining high fidelity. Notably, at the same compression ratio, SweetTokenizer significantly outperforms all baselines, achieving an rFVD of 18.74 on UCF-101, 57.1% lower than OmniTokenizer’s 42.35, and an rFVD of 7.51 on K-600, 71.0% lower than OmniTokenizer’s 25.97.
The generative capability of SweetTokenizer is evaluated on UCF-101 in a class-conditional generation task. SweetTokenizer is used to extract quantized spatial tokens ($\hat{\Omega}_{s}$) and temporal tokens ($\hat{\Omega}_{t}$) from UCF videos. These tokens are then concatenated to form training sequences for a 24-layer autoregressive (AR) transformer [49], following the same generation protocol as OmniTokenizer. As shown in Table 2, SweetTokenizer achieves a significant performance improvement, with a gFVD score of 128, 32.9% lower than OmniTokenizer’s 191 gFVD. This improvement is attributed to SweetTokenizer’s effective token compression, which substantially reduces the training complexity for downstream autoregressive models.
The visualization results are presented in Figure 5. To ensure a fair comparison, we select generated videos with similar appearances, as the generation process inherently involves randomness. The results demonstrate significantly improved detail, such as clearer human facial features and finer table textures. Additionally, SweetTokenizer effectively preserves temporal consistency, even under large motion scenarios. Further quantitative and qualitative results on video datasets are provided in the supplementary.
4.3 Image Reconstruction
We evaluate the image tokenization performance of SweetTokenizer on ImageNet by fine-tuning the spatial branch $CQAE_{s}$ with a pixel-level loss during Stage 3. As shown in Table 3, we compare SweetTokenizer with recent methods under various token compression settings. With 256 spatial tokens, SweetTokenizer outperforms TiTok [56] by 27.8%, reducing rFID from 1.01 to 0.73. When using 1,024 spatial tokens, SweetTokenizer achieves a significant improvement over both VQ-based and non-VQ-based methods (marked $\diamond$), achieving an rFID of 0.37, which surpasses LlammaGen-8 [35] by 37.3%. More quantitative and qualitative results on image reconstruction and generation tasks are shown in the supplementary.
4.4 Ablation Studies
Spatial-Temporal Decoupling.
We demonstrate that naively flattening video tokens into a 1D sequence is infeasible, as shown in Table 4. Training SweetTokenizer without decoupling results in significantly degraded performance. We attribute this degradation to two factors: (1) the flattening operation discards substantial consecutive temporal information, and (2) without decoupling, the model cannot leverage the crucial spatial pretraining from large-scale image datasets, as further discussed in the next paragraph.
Progressive Training Schedule.
This paragraph highlights the advantages of our progressive training schedule for the model, evaluated on the UCF-101 dataset. The results, presented in Table 4, demonstrate the critical role of each training stage. First, we underscore the importance of spatial training: omitting Stage 1 results in significantly poor rFVD scores of 269.34 and 501.12 for the combinations of “Stage 2 + Stage 3” and “Stage 2” alone, respectively. Secondly, the proxy code finetuning strategy employed in Stage 3 has a substantial impact on overall performance, improving the rFVD score from 155.92 with “Stage 1 + Stage 2” to 44.35 with the full sequence of “Stage 1 + Stage 2 + Stage 3”. These findings highlight the necessity of our complete progressive learning schedule, as each stage contributes to enhancing the model performance.
Architecture & Training Schedule rFVD $\downarrow$
Stage2 w/o decoupling 892.70
Stage2 501.12
Stage2 $+$ Stage3 269.34
Stage1 $+$ Stage2 155.92
Stage1 $+$ Stage2 $+$ Stage3 (SweetTokenizer) 44.35
Table 4: Ablation study of spatial-temporal decoupling and progessive training schedule.
UCF-101 rFVD $\downarrow$
Baseline (w/o LLC) 67.94
$+$ Spatial LLC 55.11
$+$ Spatial & Temporal LLC (SweetTokenizer) 44.35
$+$ CLIP [31]-based LLC 45.56
$+$ Qwen [1]-based LLC (SweetTokenizer) 44.35
ImageNet rFID $\downarrow$
Baseline (w/o LLC) 1.26
$+$ Spatial LLC (SweetTokenizer) 0.73
Table 5: Ablation study of language-based latent codebook (LLC) settings on UCF-101 and ImageNet.
Language-based Latent Codebook.
We evaluate the impact of our architecture design on model performance, focusing on the effects of the spatial-temporal language-based latent codebooks (LLC) on the ImageNet and UCF-101 datasets. As illustrated in Table 5, the integration of the spatial language codebook consistently enhances performance across both video and image modalities, improving the rFVD score from 67.94 to 55.11 and the rFID score from 1.26 to 0.73. Moreover, our temporal language codebook significantly benefits video reconstruction tasks, further reducing the rFVD score from 55.11 to 44.35. This underscores the importance of our unique design for video modalities. We attribute these improvements to the robust semantic features provided by language-based embeddings and their inherent alignment between visual and text spaces. Additionally, we compare different types of language-based embeddings, such as using CLIP [31] embeddings in place of Qwen-2.5B [1] embeddings. The experiments indicate that while CLIP embeddings yield satisfactory results, employing a more powerful pre-trained language model leads to superior performance.
4.5 Visual Semantic Comprehension
Few-Shot Visual Classification.
To evaluate the semantic capabilities of SweetTokenizer, we conducted experiments on few-shot image classification and video action recognition tasks. In both experiments, we initially extracted visual tokens using SweetTokenizer and transformed them into natural language words via our LLM codebook. Subsequently, we employed CLIP to compute the similarity between the visual inputs and text embeddings. The top 21 tokens with the highest similarity were selected to form a prompt for prediction using the Qwen LLM. For the image classification task, we adhered to the V2L protocol [59], comparing SweetTokenizer against other language-based visual tokenizers SPAE [54] and V2L. In the video action recognition task, we used ARN [58] and HF-AR [25] as baselines. The results in Table 6 indicate that SweetTokenizer achieved an accuracy of 90.8% on the miniImageNet dataset, surpassing SPAE 89.9% and V2L 87.6%. On the UCF-101 dataset, SweetTokenizer attain an average accuracy of 90.1%, outperforming ARN 83.1% and HF-AR 86.4%. These findings demonstrate SweetTokenizer robust semantic understanding and superior performance in both image and video tasks.
Methods ImageNet UCF-101
K-way-N-shot 2-1 2-3 2-5 Avg 5-5
SPAE [54] 84.8 92.5 92.6 89.9 -
V2L [59] 76.3 91.2 95.3 87.6 -
ARN[58] - - - - 83.1
HF-AR [25] - - - - 86.4
SweetTokenizer 86.8 90.5 95.2 90.8 90.1
Table 6: Few-shot visual classification accuracy ($\uparrow$), evaluated on both image and video modality.
5 Conclusions
We present SweetTokenizer, an efficient visual tokenization framework that compresses spatial and temporal information through the cross-attention query autoencoder. Combined with language-based latent spaces and progressive training schedule, SweetTokenizer reduces token count by 4$\times$ for video data while maintaining high reconstruction fidelity. Our approach offers a compact representation of visual data, making it well-suited for downstream tasks such as visual generation, visual recognition, marking a significant step in efficient visual tokenization.
Supplementary Material
Notations Explanations
$CQAE_{s,t}$ Cross-attention query autoencoder
$\mathcal{E}$ Encoder
$\mathcal{D}$ Decoder
$\mathcal{Q}$ Quantizer
$p_{t,h,w}$ Downsample ratio
$\Omega_{s,t,m}$ Continuous latent query tokens
$z_{s,t,m}$ Continuous latent query token
$\hat{\Omega}_{s,t}$ Quantized latent query tokens
$\hat{z}_{s,t}$ Quantized latent query token
$v_{s,t}$ Continuous visual feature
$\hat{v}_{s,t}$ Discrete visual feature
$h(\cdot)$ Hidden unit
$\Delta v$ Difference of consecutive features
$C_{adj,noun,adverb,verb}$ Codebook text embeddings
$c_{adj,noun,adverb,verb}$ Codebook text embedding
$\mathcal{P}$ Projector network
$sg(\cdot)$ Stop gradient operation
$\Psi_{e,d}(\cdot)$ Pretrained encoder & decoder
$\mathcal{L}_{c}$ Cross-Entropy loss
$\mathcal{L}_{vq}$ Commitment loss
$\mathcal{L}_{m}$ Mean-squared loss
$\mathcal{L}_{g}$ Gan loss
$\mathcal{L}_{p}$ Perceptual loss
Table 7: Explanations for the notations in the main paper.
6 Experimental Settings
6.1 Model Implementation Details
Visual Tokenizer.
The tokenizer is composed of an encoder $\mathcal{E}$, decoder $\mathcal{D}$, and latent quantizer $\mathcal{Q}$. The tokenizer takes a video clip of 17 consecutive frames with a resolution of 256 $\times$ 256 with the elements normalized to $[-0.5,0.5]$ as input. Then the video clip will be patchified to a resolution of 1 $\times$ 32 $\times$ 32 spatial feature and 4 $\times$ 32 $\times$ 32 temporal feature as illustrated in the main paper. The encoder $\mathcal{E}$ and decoder $\mathcal{D}$ in our tokenizer are both composed of 8 $CQAE_{s}$ modules and 4 $CQAE_{t}$ modules with 512 hidden states and 8 attention heads. Each modules consists of self-attention, feed-forward and cross-attention layers. Before the attention computation, the visual features will be reshaped into $[(BT)\times(HW)\times D]$ and $[(BHW)\times T\times D]$ for $CQAE_{s}$ and $CQAE_{t}$ modules, respectively. The encoder $\mathcal{E}$ generates 256 spatial and 1024 temporal continuous latent tokens. These tokens are then passed to the quantizer $\mathcal{Q}$, which produces the quantized spatial and temporal latent tokens. The quantizer $\mathcal{Q}$ is composed of a spatial and temporal codebook and a GCN with two hidden layers with hidden dimension of 512 as the projector network $\mathcal{P}$. To improve the training stability of the visual tokenizer, we adopt exponential moving average (EMA) updates with weight of 0.999 following [56].
Language Model.
We utilize VideoGPT following [45] as the default large language model for the video generative pre-training. All settings follow the protocol of [45].
Tokenizer #Params #Tokens FID $\downarrow$
VQVAE-2 [32] 13.5B 1024 31.11
VQGAN [11] 1.4B 1024 74.3
RQ-Transformer [26] 821M 1024 13.11
ViT-VQGAN [46] 650M 1024 8.81
OmniTokenizer [45] 650M 1024 7.45
SweetTokenizer 650M 256 5.48
Table 8: Comparison of class-conditional generation results on ImageNet-1K. Each image is of resolution of 256 $\times$ 256. We only compare autoregressive generators (AR) as baselines. “#Params” refers the parameter number of generator.
Configuration Language Model Tokenizer
Stage 1 Stage 2 Stage 3
LLM init VideoGPT - - -
Optimizer AdamW AdamW AdamW AdamW
Optimizer Hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.96$ $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1e^{-8}$ $\beta_{1}=0.9$, $\beta_{2}=0.999$
Batch size per GPU 4 32 8 12
Peak learning rate $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
Discriminator peak learning rate - - - $1e^{-4}$
Learning rate schedule Cosine Cosine Cosine Cosine
Training steps 1000K 1000K 500K 500K
Discriminator start steps - - - 20K
Warm-up steps 10K 10K 10K 10K
Weight decay 0.03 $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
Numerical precision float16 float16 float16 bfloat16
Table 9: The detailed training hyperparameters of SweetTokenizer.
6.2 Training Datasets
UCF-101.
UCF-101 is a large-scale action recognition dataset consisting of 13,320 videos with 9537 for training and 3783 for testing across 101 action categories. The dataset includes videos with significant variations in camera motion, object appearance, scale, viewpoint, cluttered backgrounds, and lighting conditions, making it one of the most challenging datasets for action recognition.
Kinetic-600.
Kinetics-600 is a large-scale action recognition dataset containing approximately 480K videos across 600 action categories. The dataset is split into 390K training, 30K validation, and 60K test videos. Each video is a 10-second clip extracted from raw YouTube footage, focusing on key action moments.
ImageNet-1K.
ImageNet-1K is a widely used subset of the larger ImageNet dataset, specifically designed for image classification tasks. It contains 1.2 million labeled images across 1,000 distinct categories, ranging from animals and plants to everyday objects and scenes. Each category in ImageNet-1K includes a set of training images, along with separate validation and test sets for model evaluation. The dataset is widely used for researchs in computer vision.
6.3 Notations
The meaning of our notations appeared in the main paper are explained in Table 7.
6.4 Training Settings
The detailed training hyper-parameter settings for SweetTokenizer are reported in Table 9.
7 Additional Results
7.1 Generation results for ImageNet
Following the OmniTokenizer protocol [45], we train a VideoGPT-based generator on ImageNet-1K. As shown in Table 8, SweetTokenizer consistently outperforms all baselines in terms of FID. With fewer tokens, SweetTokenizer achieves 5.48 FID, outperforming OmniTokenizer by 26.4 %. These results align with findings on UCF-101, reinforcing that fewer tokens lead to better generation performance.
7.2 More Visualizations
Fig 8 and Fig 7 visualize the reconstruction results for the UCF-101 and K-600 datasets. The pixel-level differences between ground truth and model are shown, with brighter areas indicating greater disparity and darker areas reflecting consistency. As shown, SweetTokenizer exhibits fewer reconstruction differences compared to OmniTokenizer, demonstrating its superior performance.
Fig 8 and Fig 9 visualize the reconstruction and generation results of SweetTokenizer on ImageNet-1K. For reconstruction, differences between models are highlighted in red blocks, with details shown in green blocks. Clearly, SweetTokenizer outperforms all baselines by a significant margin.
Finally, we visualize the words from our LLM codebook in Fig 10, based on few-shot video action recognition tasks on the UCF-101 dataset. We use adjectives, nouns, adverbs, and verbs as prompts to Qwen LLM for action prediction. Green and orange indicate meaningful words, while red marks meaningless ones. The visualization shows that correct verb words consistently lead to accurate predictions, even when other words are irrelevant, highlighting the importance of our adverb and verb LLM codebook modules for video action recognition.
Figure 6: Comparison of the reconstruction results of OmniTokenizer and SweetTokenizer on UCF-101 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
Figure 7: Comparison of the reconstruction results of OmniTokenizer and SweetTokenizer on K-600 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
Figure 8: Comparison of the reconstruction results of TiTok, OmniTokenizer, and SweetTokenizer on ImageNet-1K dataset. Differences are selected by the red blocks and highlighted in the grean blocks.
Figure 9: Class-conditional image generation results using SweetTokenizer, on ImageNet-1k dataset.
Figure 10: Semantic words visualization for UCF-101. The visualization is based on few shot video action recognition tasks.
8 Limitations
Our tokenizer is not suitable for tasks requiring precise semantic understanding, like VQA, because the LLM codebook is trained in an unsupervised manner. Without additional constraints, such as contrastive learning between image features from Qwen-VLM and text embeddings in our codebook, aligning the image and text domains is challenging. A promising direction for future work is to enhance SweetTokenizer into a semantically strong tokenizer by contrastive learning.

### v3

@@
-SweetTokenizer: Semantic-Aware Spatial-Temporal Tokenizer for Compact Visual Discretization
+SweetTok: Semantic-Aware Spatial-Temporal Tokenizer for Compact Video Discretization
 Abstract
-This paper presents the Semantic-aWarE spatial-tEmporal Tokenizer (SweetTokenizer), a compact yet effective discretization approach for vision data.
-Our goal is to boost tokenizers’ compression ratio while maintaining reconstruction fidelity in the VQ-VAE paradigm.
-Firstly, to obtain compact latent representations, we decouple images or videos into spatial-temporal dimensions, translating visual information into learnable querying spatial and temporal tokens through a Cross-attention Query AutoEncoder (CQAE).
-Secondly, to complement visual information during compression, we quantize these tokens via a specialized codebook derived from off-the-shelf LLM embeddings to leverage the rich semantics from language modality.
-Finally, to enhance training stability and convergence, we also introduce a curriculum learning strategy, which proves critical for effective discrete visual representation learning.
-SweetTokenizer achieves comparable video reconstruction fidelity with only 25% of the tokens used in previous state-of-the-art video tokenizers, and boost video generation results by 32.9% w.r.t gFVD.
-When using the same token number, we significantly improves video and image reconstruction results by 57.1% w.r.t rFVD on UCF-101 and 37.2% w.r.t rFID on ImageNet-1K.
-Additionally, the compressed tokens are imbued with semantic information, enabling few-shot recognition capabilities powered by LLMs in downstream applications.
+This paper presents the Semantic-aWarE spatial-tEmporal Tokenizer (SweetTok), a novel video tokenizer to overcome the limitations in current video tokenization methods for compacted yet effective discretization.
+Unlike previous approaches that process flattened local visual patches via direct discretization or adaptive query tokenization, SweetTok proposes a decoupling framework, compressing visual inputs through distinct spatial and temporal queries via Decoupled Query AutoEncoder (DQAE).
+This design allows SweetTok to efficiently compress video token count while achieving superior fidelity by capturing essential information across spatial and temporal dimensions.
+Furthermore, we design a Motion-enhanced Language Codebook (MLC) tailored for spatial and temporal compression to address the differences in semantic representation between appearance and motion information.
+SweetTok significantly improves video reconstruction results by 42.8% w.r.t rFVD on UCF-101 dataset.
+With a better token compression strategy, it also boosts downstream video generation results by 15.1% w.r.t gFVD.
+Additionally, the compressed decoupled tokens are imbued with semantic information, enabling few-shot recognition capabilities powered by LLMs in downstream applications.
 1 Introduction
 Figure 1: Illustration of our framework.
-We build a compact visual latent space by reducing token count and leveraging semantic text embedding.
+We build a compact visual latent space by reducing token count in a decoupled style and leveraging motion-enhanced semantic text embedding.
 The encoded tokens can be applied to downstream tasks, such as generation and understanding.
-Visual tokenizers [41, 6, 45, 43, 12, 53, 55] are emerging as essential components in the field of modern computer vision models, particularly in the generation [12, 55, 50] and understanding [44, 18, 36, 28, 47] of vision data.
+Visual tokenizers [36, 5, 41, 38, 11, 48, 50, 39] are emerging as essential components in the field of modern computer vision models, particularly in the generation [11, 50, 45] and understanding [40, 14, 32, 23, 42] of vision data.
 These tools convert visual inputs into discrete tokens, capturing essential temporal and spatial features that facilitate advanced analysis by formulating visual-related tasks as a token prediction process.
 Compression ratio and reconstruction fidelity are vital criteria for evaluating a tokenizer.
-However, recent visual tokenizers, especially video tokenizers [45, 43, 55] typically retain a low compression ratio.
+Recent visual tokenizers, especially video tokenizers [41, 38, 50] typically retain a low compression ratio.
 This is because visual tokens are usually derived from 2D patches [9] or 3D tubes [41, 11] which preserve location relationships (e.g., each token corresponds to a specific region of input [51]), leading to redundancy in both spatial and temporal dimensions.
-To reduce token count, we take inspiration from Q-Former [27], which compresses raw input into learnable “query tokens”.
-However, it is observed that directly flattening video tokens into sequence may lead to catastrophic temporal information loss.
-Therefore, we explore a spatial-temporal decoupled approach to perform tokenization.
-A higher compression ratio generally makes reconstruction more challenging.
-To complement visual information during compression, we follow [29, 54, 57] to introduce pretrained language embeddings as the latent codebook, leveraging the powerful semantic representation capabilities of large language models (LLM) [1, 38].
+Most recent work LARP [39] quantizes the flattened video patches through adaptive holistic queries to achieve high compression ratio.
+However, it is observed that directly flattening video tokens into sequence may lead to difficulty in learning intertwined spatial temporal information resulting in low reconstruction performance.
+Therefore, a new compression method needs to be proposed, one that takes into account the spatiotemporal properties of video.
+Another issue, meanwhile, is that a higher compression ratio typically results in a greater loss of reconstruction details.
+To complement visual information during compression, one common strategy is to introduce pretrained language embeddings as the latent codebook [24, 49, 52] , leveraging their semantic representation capabilities.
 However, previous works primarily focus on image modality, overlooking the relationships between text and motion in video domain.
-To address existing limitations, we propose SweetTokenizer – Semantic-aWarE spatial-tEmporal Tokenizer – as illustrated in Figure 1.
-First, considering the heterogeneous redundancy in static images and dynamic frames, we propose the Cross-attention Query AutoEncoder (CQAE) to compress spatial and temporal information into separate learnable queries.
-Different from previous works [56, 27], our findings indicate that coupling the compression of spatiotemporal information increases the difficulty for the decoder to learn the motion information of the same pixel across consecutive frames.
+To address existing limitations, we propose SweetTok – Semantic-aWarE spatial-tEmporal Tokenizer – as illustrated in Figure 1.
+Considering the heterogeneous redundancy in static images and dynamic frames, we propose the Decoupled Query AutoEncoder (DQAE) to compress spatial and temporal information into separate learnable queries.
+Different from previous works [51, 22, 39], our findings indicate that coupling the compression of spatiotemporal information increases the difficulty for the decoder to learn the motion information of the same pixel across consecutive frames.
 Thus, taking the decoupled spatial and temporal queries as inputs, we devise a strategy of spatial decoding followed by temporal decoding to achieve a separate reconstruction of the spatial and temporal dimensions of visual information.
-Additionally, the decoupled spatiotemporal reconstruction approach allows for pre-training with a large amount of image data, enhancing the model spatial representation.
-Second, to integrate the semantic information inherent in large language model (LLM), we design specialized codebooks tailored for spatial and temporal compression addressing the differences in semantic representation between spatial and temporal information.
+Additionally, the decoupled spatiotemporal reconstruction approach naturally allows for finetuning on image data, making our SweetTok flexible to image reconstruction task.
+Furthermore, to integrate the semantic information inherent in pre-traiend language model, we design a Motion-enhanced Language Codebook (MLC) tailored for spatial and temporal compression addressing the differences in semantic representation between spatial and temporal information.
 Specifically, we design two language-based codebooks based on the part of speech, using nouns and adjectives for spatial static information and verbs and adverbs for temporal motion information.
 By incorporating language-based codebooks, the learnable compressed queries can also be easily adapted to downstream visual understanding tasks by in-context learning of LLM.
-Third, we introduce the curriculum learning [2] mechanism by training SweetTokenizer in three stages to achieve stable convergence.
-We initially pre-train the spatial CQAE with image data, then joint training of the spatial and temporal CQAE with video data.
-The first two stages are supervised by the commitment loss of the codebook and the cross-entropy loss of the proxy code [56] from the pretrained tokenzier.
-In the last stage, we train the whole network and introduce the spatiotemporal decoder to reconstruct the original video.
-Exhaustive experiments demonstrate the effectiveness of SweetTokenizer.
-Compared with SOTA methods [45], SweetTokenizer achieves comparable rFVD on the UCF-101 [34] while utilizing only 25% of the tokens.
-Notably, it enhances the gFVD metric from 191 to 128.
-At an equivalent compression rate on UCF-101, SweetTokenizer reduces rFVD from 42 to 18 and gFVD from 191 to 154.
-On the ImageNet-1k [8], SweetTokenizer demonstrates a substantial improvement in rFID, decreasing it from 0.59 to 0.37.
+Exhaustive experiments demonstrate the effectiveness of SweetTok.
+Compared with vanilla video tokenizer without token compression (OmniTok[41]), SweetTok improves rFVD by 52.3% on UCF-101 [30] using only 25% of the tokens.
+Compared with vanilla query-based tokenizer (LARP [39]), SweetTok reduces rFVD from 35.15 to 20.46 and gFVD from 99 to 84, on UCF-101.
+By directly finetuning the decoupled spatial branch on the ImageNet-1k [7], SweetTok also demonstrates a substantial improvement in rFID, decreasing it from 0.59 to 0.37.
 In summary, our work makes the following key contributions:
-  • We introduce SweetTokenizer, a cutting-edge visual tokenizer that achieves decent reconstruction fidelity with a high compression ratio via spatial-temporal decoupling and cross-attention query autoencoder, reaching a “sweet spot” between compression and fidelity.
-  • We leverage a semantic-enhanced latent codebook to utilize the off-the-shell representation capability of LLM embeddings, which improves reconstruction quality and facilitates downstream video understanding tasks.
-  • We design a progressive training schedule following the curriculum learning mechanism, which ensures better convergence of visual tokenizers.
-  • We perform extensive experiments to verify the effectiveness of SweetTokenizer, which exhibits the state-of-the-art performance on video reconstruction, image reconstruction, and class-conditional video generation tasks, leading by a large margin of 57.1%, 37.2%, and 32.9%.
-Figure 2: Main pipeline.
-Our semantic-aware spatial-temporal tokenizer (SweetTokenizer) is trained in a vector-quantized variational autoencoder (VQ-VAE) style.
-We decouple video into spatial-temporal domains, then encode/decode video via the proposed novel cross-attention query autoencoder (CQAE), which significantly reduces the token number from 5,120 to 1,280 (Section 3.2.2).
-We further enhance our model with a language-based latent codebook, improving reconstruction fidelity and adding extra semantics (Section 3.2.3).
-A curriculum learning strategy is adopted to facilitate our training in three stages: Stage 1 and Stage 2 use loss marked purple; Stage 3 uses loss marked blue (Section 3.2.4).
-The quantized spatial-temporal latent queries $\hat{\Omega}_{s},\hat{\Omega}_{t}$ are used for downstream tasks.
+  • We introduce SweetTok, a cutting-edge video tokenizer that achieves the state-of-the-art reconstruction fidelity with a high compression ratio via spatial-temporal decoupling and decoupled query autoencoder (DQAE), reaching a “sweet spot” between compression and fidelity.
+  • We propose a motion-enhanced language codebook (MLC) to more effectively capture the action information embedded in the video modality, thereby improving reconstruction quality and supporting downstream video understanding tasks.
+  • We perform extensive experiments to verify the effectiveness of SweetTok, which exhibits the state-of-the-art performance on video reconstruction, image reconstruction, and class-conditional video generation tasks, leading by a large margin of 42.8%, 37.2%, and 15.1%.
+Figure 2: Pipeline overview.
+(a) Vanilla video tokenizers directly quantize flattened video patches.
+(b) Vanilla query-based tokenizers compress flattend video patches into adaptive queries.
+(c) SweetTok proposes decoupled query-based autoencoder (DQAE, §3.2.2).
+The spatial encoder quantizes the first frame’s patch embeddings, while the temporal encoder quantizes residual between consecutive frames.
+The spatial decoder reconstructs the first frame’s patches, replicates them $T$ times, and passes them to the temporal decoder for final information fusion and reconstruction.
+It also proposes motion-enhanced language codebook (MLC, §3.2.3) to complement reconstructed video information via action-related language semantics.
 2 Background
@@
 OmniTokenizer [41] and MAGVIT [48, 50] adopted similar transformer architecture and introduced image pre-training to improve video tokenizer.
+LARP [39], on the other hand, introduces to compress flattened video patches into adaptive holistic queries with the guidance of a pre-trained auto-regressive model.
 In this paper, we inherit the popular spatial-temporal decomposition design for video data.
-Moreover, we significantly increase the token compression ratio with a novel stacked cross-attention query autoencoder, which is inspired by Q-Former [27]’s token compression mechanism.
 2.2 Language-based Latent Codebook
@@
 Furthermore, De-Diffusion [43] directly encoded image into plain text as latent space interface and decodes back through a text-to-image (T2I) diffusion model [29].
-However, none of these works dives deeply into the codebook design for video modality.
-Therefore, we propose splitting the codebook according to the video’s spatial-temporal attribute, which uses nouns & adjectives for spatial information, and verbs & adverbs for temporal information.
-This design helps us better align visual-motion-text semantics.
+However, these studies primarily focus on the image modality.
+In this paper, we explore the design of the language codebook specifically for the video modality by splitting the codebook according to the video’s spatial-temporal attribute.
 3 Method
 3.1 Preliminary
-A typical visual vector-quantization (VQ) model [12, 45, 53, 55] contains three parts: encoder $\mathcal{E}$, decoder $\mathcal{D}$ and latent quantizer $\mathcal{Q}$.
-Take video modality as an example, given a video input $x\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ represents the temporal length and $H\times W$ denotes spatial resolution, encoder $\mathcal{E}(x)$ projects it into latent space $\mathbb{Z}\in\mathbb{R}^{N\times D}$, where $D$ is latent dimension and $N$ is token number.
-A quantizer $\mathcal{Q}$ is constructed in this latent space $\mathbb{Z}$ by querying the nearest neighbor in codebook $C\in\mathbb{R}^{L\times D}$, where $L$ is codebook size.
+A typical visual vector-quantization (VQ) model [11, 41, 48, 50] contains three parts: encoder $\mathcal{E}$, decoder $\mathcal{D}$ and latent quantizer $\mathcal{Q}$ as shown in Figure 2 (a).
+Take video modality as an example, given a video input $x\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ represents the temporal length and $H\times W$ denotes spatial resolution, encoder $\mathcal{E}(x)$ projects it into latent space $\mathbb{Z}\in\mathbb{R}^{N\times D}$, where $D$ is latent dimension and $N$ is token number, A quantizer $\mathcal{Q}$ is constructed in this latent space $\mathbb{Z}$ by querying the nearest neighbor in codebook $C\in\mathbb{R}^{L_{c}\times D}$, where $L_{c}$ is codebook size.
 Then $\mathcal{D}$ decodes latent space back to pixel space and applies self-supervised reconstruction loss:
 $\mathcal{L}_{rec}(x,\mathcal{D}(\mathcal{Q}(\mathcal{E}(x)))).$ (1)
-We aim to learn a more compact latent space $\mathbb{Z}$ while maintaining the reconstruction fidelity.
-To this end, we design a semantic-enhanced spatial temporal decoupled tokenizer (SweetTokenizer) to address the issue of token compression (Figure 2).
-The main component is Cross-attenion Query AutoEncoders (CQAE), which aggregates information from disentangled spatial-temporal dimensions into learnable queries, achieving a 4$\times$ compression ratio improvement over previous video modality tokenizers.
-To enhance latent space representation and reconstruction fidelity, we incorporate compact text embeddings and refine vocabularies by appearance/motion attributes for better semantic alignment.
-Finally, we propose a progressive training schedule to stabilize convergence.
-3.2 Spatial-Temporal Tokenization
+Unlike traditional visual VQ models, LARP [39] quantizes flattened video patches into holistic queries $\mathbf{Q}\in\mathbb{R}^{L\times D}$ as shown in Figure 2 (b), where $L$ is the adaptive query size.
+As shown in Equation (2), the encoder $\mathcal{E}$ processes concatenated flattened video patches $\mathbf{E}=flatten(x)\in\mathbb{R}^{N\times D}$ and query tokens $\mathbf{Q}\in\mathbb{R}^{L\times D}$, outputting $\mathbf{Z_{Q}}\in\mathbb{R}^{L\times D}$ for quantization.
+The decoder $\mathcal{D}$ reconstructs the video patches $\tilde{\mathbf{x}}\in\mathbb{R}^{T\times H\times W\times 3}$ from the concatenation of learnable video queries $\mathbf{E_{Q}}\in\mathbb{R}^{N\times D}$ and the query discrete embeddings $\tilde{\mathbf{Z}}_{\mathbf{Q}}$.
+$\mathbf{Z_{E}}||\mathbf{Z_{Q}}=\mathcal{E}(\mathbf{E}||\mathbf{Q}),\tilde{\mathbf{Z}}_{\mathbf{Q}}=\mathcal{Q}(\mathbf{Z_{Q}}),\tilde{x}=\mathcal{D}(\mathbf{E_{Q}}||\tilde{\mathbf{Z}}_{\mathbf{Q}}).$ (2)
+However, directly compressing video from flattened patches is challenging due to intertwined redundant temporal data and low-level spatial content, leading to sub-optimal performance.
+Thus, we designed SweetTok to balance reconstruction performance with a high compression ratio.
+Details are elaborated in Section 3.2.
+3.2 Decoupled Spatial-Temporal Tokenization
+As noted in Section 3.1, directly quantizing video data via flattened patches hampers model learning due to intertwined redundant temporal and complex spatial information.
+Thus, we propose separately quantizing spatial and temporal dimensions before combining them to reconstruct the video, following the divide-and-conquer principle.
+The decoupling strategy enables high compression while ensuring higher fidelity.
+The main pipeline is shown in Figure 2 (c).
 3.2.1 Patchify
-Given a video frame sequence $x\in\mathbb{R}^{T\times H\times W\times 3}$, we select the first frame $x_{0}$ as a reference for spatial information, the remaining $T-1$ frames $x_{1:T}$ for temporal information, following the strategy in [45].
-We apply two patch kernels with shapes $p_{h}\times p_{w}$ and $p_{t}\times p_{h}\times p_{w}$ to $x_{0}$ and $x_{1:T}$ separately, generating $v_{s}\in\mathbb{R}^{1\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ and $v_{t}\in\mathbb{R}^{\frac{T-1}{p_{t}}\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ as inputs for transformer-based autoencoder, where $v_{s}$ contains spatial information, $v_{t}$ contains temporal information, with $D$ representing the hidden embedding size.
+Given a video frame sequence $x\in\mathbb{R}^{T\times H\times W\times 3}$, we select the first frame $x_{1}$ as a reference for spatial information, the remaining $T-1$ frames $x_{2:T}$ for temporal information, following the strategy in [41].
+We apply two patch kernels $\mathcal{P}_{s},\mathcal{P}_{t}$ with shapes $p_{h}\times p_{w}$ and $p_{t}\times p_{h}\times p_{w}$ to $x_{1}$ and $x_{2:T}$ separately, generating $v_{s}\in\mathbb{R}^{1\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ and $v_{t}\in\mathbb{R}^{\frac{T-1}{p_{t}}\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ shown below:
+$v_{s}=\mathcal{P}_{s}(x_{1}),v_{t}=\mathcal{P}_{t}(x_{2:T})$ (3)
+$v_{s}$ and $v_{t}$ are inputs for transformer-based autoencoder, where $v_{s}$ contains spatial information, $v_{t}$ contains temporal information.
 In practice, for a video with 17 frames and a resolution of $256\times 256$, we set $(p_{t},p_{h},p_{w})$ to (4, 8, 8), thus patchify frames into $v_{s}$ with shape $1\times 32\times 32$ and $v_{t}$ with shape $4\times 32\times 32$.
-In Figure 2, we use $t=\frac{T-1}{p_{t}}=4$ to denote $v_{t}$’s length.
-Figure 3: Module structure of Cross-attention Query AutoEncoder (CQAE).
-Information is transported from $K,V$ head to $Q$ head through stacked CQAE transformer layers.
-The enlarged view illustrates the detailed architecture.
-3.2.2 Cross-attention Query AutoEncoder (CQAE)
-So far, the raw video has been split into patches with a fixed ratio $p_{t}\times p_{h}\times p_{w}$.
-However, many works [56, 27, 37] have demonstrated that visual information remains redundant spatially and temporally, even after tokenization.
-To select “key words” from these tokens, we adopt a strategy similar to Q-Former [27], compressing patch sequence into a fixed number of query tokens via cross-attention interactions.
-As an innovation, we recursively inject these cross-attention query modules into transformer-based autoencoder to transfer information, forming our CQAE module.
-Moreover, we show in experiments that naively flattening patches into a long 1D sequence is unsuitable for video modality since it may lose temporal motion information in consecutive frames.
-Therefore, we introduce two autoencoders, $CQAE_{s}$ and $CQAE_{t}$, along with two sets of latent query token embeddings $\Omega_{s}\in\mathbb{R}^{N_{s}\times D}$ and $\Omega_{t}\in\mathbb{R}^{N_{t}\times D}$, for summarizing spatial and temporal information, where $N_{s}$ and $N_{t}$ control the latent token number.
-We set $N_{s}$ to 256 and $N_{t}$ to 1,024, resulting in 1,280 tokens in total.
-During the encoding stage, we pass $v_{s}$ into $CQAE_{s}$ and $v_{t}$ into $CQAE_{t}$.
-Take $CQAE_{s}$ as example, as shown in Figure 3, $v_{s}$ is first reshaped from $1\times 32\times 32$ into $1\times 1024$, then passed through $K$-layer self-attention transformer to produce $K$ intermediate hidden units $[h(v_{s}^{1}),h(v_{s}^{2}),...,h(v_{s}^{K})]$.
-Meanwhile, the latent query token $\Omega_{s}$ is also passed through self-attention transformers, except that each hidden unit $h(\Omega_{s}^{i})$ is served as query tensor for cross-attention layer with $h(v_{s}^{i})$.
-Therefore, we gradually transfer spatial information from $v_{s}$ into $\Omega_{s}$ through interactions between hidden units $h(v_{s}^{i})$ and $h(\Omega_{s})$, reducing the spatial token number from 1,024 to 256.
-Similarly, $v_{t}$ is processed with $CQAE_{t}$.
-Instead of directly using $v_{t}$, we employ frame-wise residual $\Delta v_{t}$ to generate temporal hidden units $[h(\Delta v_{t}^{1}),h(\Delta v_{t}^{2}),...,h(\Delta v_{t}^{K})]$, as [18] found that frame motion residuals can be more effectively represented.
-Notably, we reshape $\Delta v_{t}$ from $4\times 32\times 32$ into $1024\times 4$, and $\Omega_{t}$ from $1024$ into $1024\times 1$ to align the transformer inputs shape.
-Thus, the temporal information is embedded into $\Omega_{t}$ from $v_{t}$ using the same cross-attention querying approach along the temporal axis.
-$\Omega_{s}$ and $\Omega_{t}$ are quantized into $\hat{\Omega}_{s}$ and $\hat{\Omega}_{t}$, which will be explained in Section 3.2.3 later.
-During the decoding stage, we use mask tokens $\Omega_{m}\in\mathbb{R}^{1\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ to receive information from $\hat{\Omega}_{s}$ and $\hat{\Omega}_{t}$, similar to the mask token prediction tasks in [9, 14, 37, 56].
-As shown in Figure 3, yellow blocks denote that the tensor contains spatial visual information, while grey blocks indicate that the tensor is information-free.
-We randomly initialize these mask tokens $\Omega_{m}$ and use them as a “canvas” to reconstruct the raw inputs by receiving visual information from $\hat{\Omega}_{s}$.
-The decoding stage consists of two steps.
-Firstly, the mask token $\Omega_{m}$ starts with shape $32\times 32$, which is reshaped into $1\times 1024$ to interact spatially with $\hat{\Omega}_{s}$.
-Each decoder layer has the same structure as the encoder layer, except that the query is $h(\Omega_{m}^{i})$ and the key/value is $h(\hat{\Omega}_{s}^{i})$.
-Afterwards, $\Omega_{m}$ is expanded from $1\times 1024$ into $5\times 1024$, in order to interact temporally with $\hat{\Omega}_{t}$.
-The final output $\Omega_{m}$ is a mixture of yellow and green block as illustrated in Figure 2, indicating that it simultaneously contains spatial and temporal information.
-3.2.3 Language-based Latent Codebook
-Previous works [29, 54, 57] have shown that text representations can enhance image VQ-VAEs , as the text provides additional semantic information from pre-trained language models.
-However, previous works mainly focus on the relationship between static image appearance and text semantics.
-Our experiment in Table 5 shows this is insufficient for video data, as static and motion information are typically embedded in different subsets of vocabularies.
-To address this, we construct two separate codebooks, the spatial quantization codebook and the temporal quantization codebook, according to vocabulary attributes, as illustrated in Figure 2.
-We first extract candidate vocabularies from video captions, which is obtained by Qwen-2.5B [1] instruct model.
-We filter out words that occur with low frequency.
-Afterward, we extract Qwen-2.5B [1] text embedding of these vocabularies to fill in the columns of our codebook $C\in\mathbb{R}^{L\times D}$ and split it into four subsets: nouns, adjectives, verbs, and adverbs.
-Note that $C_{noun},C_{adj}$ are for spatial latent tokens $\Omega_{s}$ while $C_{verb},C_{adv}$ are for temporal latent tokens $\Omega_{t}$.
-Given two encoded continuous latent vectors: $z_{s}\in\Omega_{s}$ and $z_{t}\in\Omega_{t}$, $z_{s}$ is passed through spatial quantization codebook, and $z_{t}$ is passed through temporal quantization codebook.
+We use $t=\frac{T-1}{p_{t}}=4$ to denote $v_{t}$’s length.
+3.2.2 Decoupled Query AutoEncoder (DQAE)
+To decouple spatial and temporal dimensions, we need to compress video patches separately along each dimension.
+Inspired by [51, 39] and Q-Former [22], we compress each dimension into an adaptive query tokens via cross-attention interactions.
+As an innovation, we recursively inject these cross-attention query modules into transformer-based autoencoder to transfer information, forming our DQAE module shown in the right gray box of Figure 2 (c).
+For the rest of our paper, we use $\mathcal{E}_{DQAE}$ and $\mathcal{D}_{DQAE}$ as the encoder and decoder of the $DQAE$ module for simplicity.
+Spatial Tokenization.
+We observe that for most video parts, the first frame holds the most spatial information, so we use $v_{s}$ as the input to $\mathcal{E}_{DQAE_{s}}$ for quantization shown below:
+$\mathbf{Z}_{\mathbf{Q_{s}}}=\mathcal{E}_{DQAE_{s}}(\mathbf{Q_{s}},v_{s}),\tilde{\mathbf{Z}}_{\mathbf{Q_{s}}}=\mathcal{Q}_{MLC}(\mathbf{Z}_{\mathbf{Q_{s}}}),$ (4)
+where $\mathbf{Q_{s}}\in\mathbb{R}^{L_{spatial}\times D}$ are the learnable spatial query embeddings and $\mathbf{Z_{\mathbf{Q_{s}}}}\in\mathbb{R}^{L_{spatial}\times D}$ are the output embeddings encoding information from the first frame patches $v_{s}$.
+$\tilde{\mathbf{Z}}_{\mathbf{Q_{s}}}\in\mathbb{R}^{L_{spatial}\times D}$ are the quantized spatial token embeddings and $\mathcal{Q}_{MLC}$ stands for Motion-enhanced Language Codebook (MLC) quantizer which will be elaborated in the following section.
+After quantization, we inject our informative spatial token embeddings $\tilde{\mathbf{Z}}_{\mathbf{Q_{s}}}$ into a learnable spatial patch queries $\mathbf{Q}_{v_{s}}$ through decoder $\mathcal{D}_{DQAE_{s}}$ as below:
+$\tilde{v}_{s}=\mathcal{D}_{DQAE_{s}}(\mathbf{Q}_{v_{s}},\tilde{\mathbf{Z}}_{\mathbf{Q_{s}}}),$ (5)
+where $\tilde{v}_{s}$ are the reconstructed first frame video patches.
+The temporal component which will be stated in the following, combined with $\tilde{v}_{s}$ is used for the final video reconstruction.
+We set $L_{spatial}=256$ in real implementation.
+Temporal Tokenization.
+It is observed that for video data, there is much redundancy along temporal dimension.
+It motivates us to employ frame-wise residual $\Delta v_{t}=(\Delta v_{t}^{1},\Delta v_{t}^{2},...,\Delta v_{t}^{k})_{k=\frac{H}{p_{h}}\times\frac{W}{p_{w}}}$, where $\Delta v_{t}^{i}=v_{s}^{i}-v_{t}^{i}$, for tokenization.
+We use the first frame of the video for frame-wise residual because the spatial tokenization phase reconsturct it.
+Then, the residual $\Delta v=(\Delta v_{1},\Delta v_{2},...,\Delta v_{t})_{t=\frac{T-1}{p_{t}}}$ is input to $\mathcal{E}_{DQAE_{t}}$ for temporal compression, as shown below:
+$\mathbf{Z}_{\mathbf{Q_{t}}}=\mathcal{E}_{DQAE_{t}}(\mathbf{Q_{t}},\Delta v),\tilde{\mathbf{Z}}_{\mathbf{Q_{t}}}=\mathcal{Q}_{MLC}(\mathbf{Z}_{\mathbf{Q_{t}}}),$ (6)
+where $\mathbf{Q_{t}}\in\mathbb{R}^{L_{temporal}\times D}$ are the learnable temporal query embeddings and $\mathbf{Z_{\mathbf{Q_{t}}}}\in\mathbb{R}^{L_{temporal}\times D}$ are the output embeddings encoding information of the frame-wise residual.
+$\tilde{\mathbf{Z}}_{\mathbf{Q_{t}}}\in\mathbb{R}^{L_{temporal}\times D}$ are the quantized frame-wise temporal residual token embeddings.
+In practical implementation, $L_{temporal}=1024$.
+After reconstructing the first frame patches, we recover the entire video patches by combining $\tilde{v}_{s}$ with the frame-wise quantized residual $\tilde{\mathbf{Z}}_{\mathbf{Q_{t}}}$.
+We tiled $\tilde{v}_{s}$ for $t$ times and sent it to $\mathcal{D}_{DQAE_{t}}$ shown below:
+$\tilde{v}=\mathcal{D}_{DQAE_{t}}([\tilde{v}_{s}||\cdots||\tilde{v}_{s}],\tilde{\mathbf{Z}}_{\mathbf{Q_{t}}}),$ (7)
+where $\tilde{v}$ is the reconstructed video patches.
+The vectors $\tilde{v}$ reside in the latent space, necessitating the use of a “pixel decoder” $\mathcal{D}_{pixel}$ to reconstruct the video data, as illustrated below:
+$\tilde{x}=\mathcal{D}_{pixel}(\tilde{v}).$ (8)
+The whole $DQAE$ is supervised by the reconstruction loss $\mathcal{L}_{rec}$ containing a $L_{2}$ loss, a LPIPS perception loss: $\mathcal{L}_{Lpips}$, a quantizer loss: $\mathcal{L}_{vq}$ and a GAN loss: $\mathcal{L}_{g}$ following the principle [10].
+3.2.3 Motion-enhanced Language Codebook (MLC)
+To mitigate information loss during compression, we introduce an language codebook (LC) quantizer to enhance semantic richness.
+The Previous works [24, 49] have shown that text representations can enhance image VQ-VAEs , as the text provides additional semantic information from pre-trained language models.
+However, previous works mainly focus on the relationship between static image appearance and text semantics [52].
+Our experiment in Table 5 shows is insufficient for video data.
+To address this, we propose a Motion-enhanced Language Codebook (MLC), where the video motion information is enhanced via action-related vocabularies.
+Specifically, we split the dictionary into four subsets: nouns, adjectives, verbs, and adverbs.
+Intuitively, static and appearance information is typically embedded in nouns and adjectives, while motion information is generally embedded in verbs and adverbs.
+Therefore, we choose nouns and adjectives for spatial query tokens $\mathbf{Q_{s}}$, and verbs and adverbs for temporal query tokens $\mathbf{Q_{t}}$.
+Figure 3 also shows that the encoded latent words by SweetTok capture semantic meanings related to both visual appearance and motion.
+As for details, we first extract candidate vocabularies of the whole dataset from video captions.
+Afterward, we extract CLIP [27] text embedding of these vocabularies to fill in the columns of our codebook $C\in\mathbb{R}^{L\times D}$.
+We utilize a graph convolution network $\mathcal{F}$ to project CLIP embeddings [27] into the visual latent space.
+Graph edges are constructed when a pair of “spatial-spatial”, “spatial-temporal” or “temporal-temporal” words co-occur within a 5-token window in the current video caption.
+Given two encoded continuous latent vectors: $z_{s}\in\mathbf{Q_{s}}$ and $z_{t}\in\mathbf{Q_{t}}$, $z_{s}$ is passed through spatial quantization codebook, and $z_{t}$ is passed through temporal quantization codebook.
 The quantized $\hat{z}_{s}$ and $\hat{z}_{t}$ are obtained by nearest neighbor searching:
 $z_{s},z_{t}=\mathcal{E}(x),$ (9)
-$\hat{z}_{s}=\mathcal{P}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{noun}\cup C_{adj}}||z_{s}-\mathcal{P}(c_{i})||,$ (3)
-$\hat{z}_{t}=\mathcal{P}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{verb}\cup C_{adv}}||z_{t}-\mathcal{P}(c_{i})||.$ (4)
-To maintain text semantic information, we freeze the codebook and use a tiny projector network $\mathcal{P}$ to map from text space into visual space, following the strategy in [57].
-Finally, the gradient is passed to the encoder via vector-quantization commitment loss proposed in [41], a common method to approximate differentiability:
+$\hat{z}_{s}=\mathcal{F}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{noun}\cup C_{adj}}||z_{s}-\mathcal{F}(c_{i})||,$ (10)
+$\hat{z}_{t}=\mathcal{F}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{verb}\cup C_{adv}}||z_{t}-\mathcal{F}(c_{i})||.$ (11)
+Finally, the gradient is passed to the encoder via vector-quantization commitment loss proposed in [36], a common method to approximate differentiability ($sg[\cdot]$ stands for stop-gradient operator):
 $\mathcal{L}_{vq}=||sg[z_{s}]-\mathcal{Q}(z_{s})||^{2}+||z_{s}-sg[\mathcal{Q}(z_{s})]||^{2}$ (12)
 $+||sg[z_{t}]-\mathcal{Q}(z_{t})||^{2}+||z_{t}-sg[\mathcal{Q}(z_{t})]||^{2}$
-where $sg[\cdot]$ is stop-gradient operator. Figure 4 also shows that the encoded latent words by SweetTokenizer capture semantic meanings related to both visual appearance and motion.
 Figure 3: The semantics of spatial-temporal “words”.
 The attention weights of the last encoder’s cross-attention layer are visualized via heatmap, showing the visual regions corresponding to the related latent words.
-3.2.4 Progressive Training Schedule
-To ensure better convergence, we follow the philosophy of curriculum learning [2], thus designing a progressive training schedule comprising three stages.
-We adopt the proxy codes strategy following [56] to leverage pre-trained visual priors.
-Assuming $\Psi_{e}(\cdot)$ is a pre-trained visual tokenizer and $\Psi_{d}(\cdot)$ is a pre-trained pixel decoder.
-$\Psi_{e}(\cdot)$ maps visual inputs $x$ into well-trained integer token indices $\Psi_{e}(x)$, which serves as a proxy code classification label of our CQAE model, similar to the knowledge distilling approach [16].
-The motivation is that we could bypass the early oscillatory training phase of GAN loss [13] as demonstrated in [56, 11], which allows us to apply it later once the training becomes stable.
-In practice, we use OmniTokenizer [45]’s encoder and decoder as $\Psi_{e}(\cdot)$ and $\Psi_{d}(\cdot)$.
-First, we train $CQAE_{s}$ on spatial patches $v_{s}$ using image data only as pre-training in Stage 1, with the supervision of proxy code:
-$\mathcal{L}_{stage1}=\mathcal{L}_{c}(CQAE_{s}(v_{s}),\Psi_{e}(x))+\mathcal{L}_{vq},$ (6)
-where $\mathcal{L}_{c}$ denotes cross entropy loss and $\mathcal{L}_{vq}$ denotes codebook commitment loss in Eq 5.
-Note that SweetTokenizer adopts spatial-temporal disentanglement, so the model weights of $CQAE_{s}$ can be later adapted to image-level tasks via finetuning.
-Experiments in Table 3 show that our spatial branch already achieves satisfactory image reconstruction results.
-Second, we jointly train $CQAE_{s},CQAE_{t}$ on $v_{s},v_{t}$ using video data in Stage 2. Since $CQAE_{s}$ is well-initialized, $CQAE_{t}$ can focus on learning temporal motions:
-$\mathcal{L}_{stage2}=\mathcal{L}_{c}(CQAE_{s}(v_{s}),\Psi_{e}(x))+$ (7)
-$\quad\quad\quad\quad\mathcal{L}_{c}(CQAE_{t}(v_{t}),\Psi_{e}(x))+\mathcal{L}_{vq}.$
-Finally, in Stage 3, we combine the pre-trained pixel decoder $\Psi_{d}(\cdot)$ together with our CQAE and finetune them by pixel-level losses in Stage 3 to boost final performance:
-$\mathcal{L}_{stage3}=\mathcal{L}_{(m,g,p)}(\Psi_{d}(CQAE(v)),x),$ (8)
-where $\mathcal{L}_{m}$ denotes mean-squared error loss, $\mathcal{L}_{g}$ denotes GAN loss [13] and $\mathcal{L}_{p}$ denotes perceptual loss [20].
 4 Experiments
 4.1 Experiments Settings
-Tokenizer #Tokens rFVD $\downarrow$
-UCF-101 K-600
-MaskGIT [6] 4352 240 202
-VQGAN [11] 4352 299 270
-TATS [12] 4096 162 -
-MAGVIT [53] 4096 58 -
-OmniTokenizer [45] 5120 42 26
-SweetTokenizer 5120 18 8
-SweetTokenizer 256 + 1024 44 28
-Table 1: Video reconstruction FVD on the UCF-101 and K-600 dataset, using a frame resolution 256 × 256. $\ast$” denotes training SweetTokenizer without token compression.
-Tokenizer #Params #Tokens gFVD $\downarrow$
-CogVideo [17] 9.4B 6800 626
-TATS [12] 321M 4096 332
-Video-LaVIT [18] 7B 512 280
-OmniTokenizer [45] 650M 5120 191
-SweetTokenizer 650M 5120 154
-SweetTokenizer 650M 256 + 1024 128
-Table 2: Comparison of class-conditional generation results on UCF-101.
-Each video is composed of 17 frames with a resolution of 256 $\times$ 256.
-We only compare autoregressive generators (AR) as baselines.
-“#Params” refers the parameter number of generator.
 Dataset.
-We evaluate the tokenization performance of SweetTokenizer on image and video datasets, including ImageNet [8], UCF-101 [34], and Kinetics-600 [21, 5].
-Following [45], all images and video frames are resized to 256 × 256 resolution for experiments.
-The semantic capabilities of SweetTokenizer are tested through few-shot image classification on Real-Name Open-Ended miniImageNet [39] and few-shot video action recognition on UCF-101, as described in [58].
+We evaluate the tokenization performance of SweetTok on video datasets, including UCF-101 [30], and Kinetics-600 [16, 4].
+Following [41], all video frames are resized to 256 $\times$ 256 resolution for experiments.
+Note that some of the previous works use a resolution of 128 $\times$ 128, which cannot be directly compared to our work due to the difference in task difficulty.
+However, we still include these results in the table and highlight them in gray.
+Moreover, we fine-tune SweetTok’s spatial component on ImageNet [7] to obtain a strong image tokenizer.
+The semantic capabilities of SweetTok are tested through few-shot image classification on Real-Name Open-Ended miniImageNet [34] and few-shot video action recognition on UCF-101, as described in [53].
 Evaluation Metrics.
@@
 For image reconstruction, we categorize recent methods by the number of compressed tokens, with each group assessed using the Frechet Inception Distance (FID) [12].
+OmniTok [41]
+LARP [39]
+SweetTok (ours)
+Figure 4: Video reconstruction result on UCF-101 dataset. We also visualize the reconstruction and GT error map, where brighter areas indicate larger errors.
+Tokenizer #Tokens #Params rFVD $\downarrow$
+Tokenizer UCF-101 K-600
+MAGVIT-V2 [50] 1280 307M 8.6 -
+LARP-L [39] 1024 193M 20 13
+MaskGIT [5] 4352 227M 240 202
+VQGAN [10] 4352 227M 299 270
+TATS [11] 4096 32M 162 -
+MAGVIT [48] 4096 158M 58 -
+OmniTok [41] 5120 82.2M 42 26
+LARP-B [39] 1024 143M 64 35
+LARP-L [39] 1024 193M 35 23
+SweetTok 5120 128M 11 8
+SweetTok 1280 128M 20 25
+Table 1: Video reconstruction FVD on the UCF-101 and K-600 dataset, using a frame resolution 256 × 256.
+$\ast$” denotes training SweetTok without token compression.
+Lines in “gray” indicate results at a resolution of 128 $\times$ 128.
+Implementation Details.
+SweetTok adopts a spatial-temporal architecture consisting of 8 spatial layers and 4 temporal layers, with both the encoder and decoder configured to a hidden dimension of 512.
+The latent space dimension is set to 256.
+For the LLM codebook quantizer, we exclude words with a frequency below 5, resulting in a selection of 5,078 nouns, 5,403 adjectives, 9,267 verbs, and 1,872 adverbs.
+This forms a spatial codebook of size 10,481 and a temporal codebook of size 11,139.
+The model is trained with a batch size of 8 for 1000K iterations.
+All training is performed on NVIDIA A100 GPUs.
+Adam [18] is employed for optimization ($\beta_{1}=0.9$ and $\beta_{2}=0.99$).
+During each stage, we use a cosine learning rate scheduler with a max learning rate of 1e-4 and a min learning rate of 1e-5, warmed up by 10K iterations.
+Tokenizer Type #Tokens #Params gFVD $\downarrow$
+Generator
+MAGVIT [48] AR 1024 306M 265
+MAGVIT-V2 [50] AR 1280 307M 109
+MAGVIT [48] MLM 1024 306M 76
+MAGVIT-V2 [50] MLM 1280 307M 58
+LARP-L [39] AR 1024 632M 57
+CogVideo [13] AR 6800 9.4B 626
+TATS [11] AR 4096 321M 332
+Video-LaVIT [14] AR 512 7B 280
+OmniTok [41] AR 5120 650M 191
+LARP-L [39] AR 1024 632M 99
+SweetTok AR 1280 650M 84
+SweetTok AR 1280 1.9B 65
+Table 2: Class-conditional video generation results on UCF-101.
+Each video is composed of 17 frames with a resolution of 256 $\times$ 256.
+“AR” and “MLM” represents autoregressive and masked-language-modeling generator.
+Lines in “gray” indicate results at a resolution of 128 $\times$ 128.
 Tokenizer #Tokens Codebook Size rFID $\downarrow$
@@
 LlamaGen-16 [31] 256 16384 2.19
-TiTok [56] 256 4096 1.03
-SweetTokenizer 256 10481 0.73
+TiTok [51] 256 4096 1.71
+TokenFlow [25] 256 4096 1.03
+SweetTok 256 10481 0.73
 ViT-VQGAN [46] 1024 8192 1.28
-OmniTokenizer [45] 1024 8192 1.11
-OmniTokenizer [45] 1024 8192 0.69
+OmniTok [41] 1024 8192 1.11
+OmniTok [41] 1024 8192 0.69
 LlamaGen-8 [31] 1024 16384 0.59
-SweetTokenizer 1024 10481 0.37
+SweetTok 1024 10481 0.37
 Table 3: Image reconstruction FID on the ImageNet dataset, using a resolution of 256 $\times$ 256.
 $\diamond$” denotes continuous latent space without quantization.
-$\ast$” denotes training SweetTokenizer without token compression.
-Implementation Details.
-SweetTokenizer adopts a spatial-temporal architecture consisting of 8 spatial layers and 4 temporal layers, with both the encoder and decoder configured to a hidden dimension of 512.
-The latent space dimension is set to 256.
-For the LLM codebook quantizer, we exclude words with a frequency below 5, resulting in a selection of 5,078 nouns, 5,403 adjectives, 9,267 verbs, and 1,872 adverbs.
-This forms a spatial codebook of size 10,481 and a temporal codebook of size 11,139.
-We utilize a graph convolution network (GCN) to project Qwen-2.5 embeddings [1] into the visual latent space.
-Graph edges are constructed when a pair of words co-occur within a 5-token window in the text.
-The model is trained with a progressive three-stage schedule: Stage 1 uses a batch size of 32 for 300K iterations, Stage 2 uses a batch size of 8 for 300K iterations, and Stage 3 uses a batch size of 4 for 500K iterations.
-All training is performed on NVIDIA A100 GPUs.
-Adam [23] is employed for optimization ($\beta_{1}=0.9$ and $\beta_{2}=0.99$).
-During each stage, we use a cosine learning rate scheduler with a max learning rate of 1e-4 and a min learning rate of 1e-5, warmed up by 10K iterations.
-OmniTokenizer [45]
-SweetTokenizer (ours)
-Figure 5: Class-conditional video generation result using OmniTokenizer [45] and our SweetTokenizer, on UCF-101 dataset.
-The action class label of each row is: “BlowingCandles”, “PlayingTabla”, “TennisSwing”, “PlayingViolin” and “HorseRiding”.
+$\ast$” denotes training SweetTok without token compression.
+OmniTok [41]
+LARP [39]
+SweetTok (ours)
+Figure 5: Class-conditional video generation result on UCF-101 dataset. The action class label of each row is: “PlayingTabla”, “TennisSwing”, and “HorseRiding”.
 4.2 Video Reconstruction & Generation
-We first evaluate the tokenization capability of SweetTokenizer on the UCF-101 and K-600 video datasets.
-As shown in Table 1, SweetTokenizer uses only 1,280 tokens (256 spatial tokens and 1,024 temporal tokens), which is four times fewer than OmniTokenizer’s 5,120 tokens.
-Despite this significant reduction, it achieves comparable performance, with rFVD scores of 44.35 on UCF-101 and 27.95 on K-600, demonstrating its effectiveness in compressing tokens while maintaining high fidelity.
-Notably, at the same compression ratio, SweetTokenizer significantly outperforms all baselines, achieving an rFVD of 18.74 on UCF-101, 57.1% lower than OmniTokenizer’s 42.35, and an rFVD of 7.51 on K-600, 71.0% lower than OmniTokenizer’s 25.97.
-The generative capability of SweetTokenizer is evaluated on UCF-101 in a class-conditional generation task.
-SweetTokenizer is used to extract quantized spatial tokens ($\hat{\Omega}_{s}$) and temporal tokens ($\hat{\Omega}_{t}$) from UCF videos.
-These tokens are then concatenated to form training sequences for a 24-layer autoregressive (AR) transformer [49], following the same generation protocol as OmniTokenizer.
-As shown in Table 2, SweetTokenizer achieves a significant performance improvement, with a gFVD score of 128, 32.9% lower than OmniTokenizer’s 191 gFVD.
-This improvement is attributed to SweetTokenizer’s effective token compression, which substantially reduces the training complexity for downstream autoregressive models.
+We first evaluate the tokenization capability of SweetTok on the UCF-101 and K-600 video datasets.
+As shown in Table 1, SweetTok uses 1,280 tokens (256 spatial tokens and 1,024 temporal tokens), which is four times fewer than OmniTok’s 5,120 tokens and comparable with LARP’s 1024 tokens.
+Despite its high compression ratio, our method achieves a 52.3% improvement on UCF-101 and competitive performance on K-600 compared with OmniTok.
+With a similar token count, SweetTok, despite its smaller model size compared to LARP-L, delivers a 42.8% improvement on UCF-101 and comparable results on K-600.
+Our SweetTok demonstrates significant performance gains, achieving 68.8% and 28.5% improvements in rFVD on UCF-101 and K-600 datasets, respectively, compared to LARP-B with similar model size.
+Notably, if we increase the token number to 5,120, SweetTok significantly outperforms all baselines, achieving an rFVD of 10.74 on UCF-101 and 7.51 on K-600.
+The generative capability of SweetTok is evaluated on UCF-101 in a class-conditional generation task.
+SweetTok is used to extract decoupled tokens from UCF videos.
+These tokens are then concatenated to form training sequences for VideoGPT [44], following the same generation protocol as OmniTok.
+As shown in Table 2, SweetTok achieves a significant performance improvement, with a gFVD score of 84, 56% lower than OmniTok’s 191 gFVD.
+This improvement is attributed to SweetTok’s effective token compression, which substantially reduces the training complexity for downstream autoregressive models.
+With equivalent token counts, SweetTok demonstrates superior performance, achieving a 15.1% improvement in gFVD (84 vs.
+LARP’s 99) at comparable generator sizes.
+Furthermore, it exhibits scaling law characteristics, with gFVD improving from 84 to 65 as the model scales up to 1.9B parameters.
 The visualization results are presented in Figure 5.
@@
 The results demonstrate significantly improved detail, such as clearer human facial features and finer table textures.
-Additionally, SweetTokenizer effectively preserves temporal consistency, even under large motion scenarios.
-Further quantitative and qualitative results on video datasets are provided in the supplementary.
+Additionally, SweetTok effectively preserves temporal consistency, even under large motion scenarios.
+Compression Method #Tokens rFVD $\downarrow$
+Vanilla Downsample 1280 227.65
+Vanilla Query-based (LARP [39]) 1024 35.15
+Decoupled Query-based (DQAE) 1280 20.46
+Table 4: Ablation study of different token count compression method for video tokenizers.
 4.3 Image Reconstruction
-We evaluate the image tokenization performance of SweetTokenizer on ImageNet by fine-tuning the spatial branch $CQAE_{s}$ with a pixel-level loss during Stage 3.
-As shown in Table 3, we compare SweetTokenizer with recent methods under various token compression settings.
-With 256 spatial tokens, SweetTokenizer outperforms TiTok [56] by 27.8%, reducing rFID from 1.01 to 0.73.
-When using 1,024 spatial tokens, SweetTokenizer achieves a significant improvement over both VQ-based and non-VQ-based methods (marked $\diamond$), achieving an rFID of 0.37, which surpasses LlammaGen-8 [35] by 37.3%.
-More quantitative and qualitative results on image reconstruction and generation tasks are shown in the supplementary.
+To demonstrate the flexibility of our decoupled query design, we show that by directly fine-tuning the spatial branch $DQAE_{s}$, SweetTok also achieves advanced performance for image reconstruction on ImageNet.
+As shown in Table 3, we compare SweetTok with recent methods under various token compression settings.
+With 256 spatial tokens, SweetTok outperforms TiTok [51] by 27.8%, reducing rFID from 1.01 to 0.73.
+When using 1,024 spatial tokens, SweetTok achieves a significant improvement over both VQ-based and non-VQ-based methods (marked $\diamond$), achieving an rFID of 0.37, which surpasses LlammaGen-8 [31] by 37.3%.
+Visualization results are shown in Figure 6, SweetTok maintains better global appearance and local details.
+GT
+TiTok [51]
+OmniTok [41]
+SweetTok
+Figure 6: Visualization of image reconstruction results.
 4.4 Ablation Studies
-Spatial-Temporal Decoupling.
-We demonstrate that naively flattening video tokens into a 1D sequence is infeasible, as shown in Table 4.
-Training SweetTokenizer without decoupling results in significantly degraded performance.
-We attribute this degradation to two factors: (1) the flattening operation discards substantial consecutive temporal information, and (2) without decoupling, the model cannot leverage the crucial spatial pretraining from large-scale image datasets, as further discussed in the next paragraph.
-Progressive Training Schedule.
-This paragraph highlights the advantages of our progressive training schedule for the model, evaluated on the UCF-101 dataset.
-The results, presented in Table 4, demonstrate the critical role of each training stage.
-First, we underscore the importance of spatial training: omitting Stage 1 results in significantly poor rFVD scores of 269.34 and 501.12 for the combinations of “Stage 2 + Stage 3” and “Stage 2” alone, respectively.
-Secondly, the proxy code finetuning strategy employed in Stage 3 has a substantial impact on overall performance, improving the rFVD score from 155.92 with “Stage 1 + Stage 2” to 44.35 with the full sequence of “Stage 1 + Stage 2 + Stage 3”.
-These findings highlight the necessity of our complete progressive learning schedule, as each stage contributes to enhancing the model performance.
-Architecture & Training Schedule rFVD $\downarrow$
-Stage2 w/o decoupling 892.70
-Stage2 501.12
-Stage2 $+$ Stage3 269.34
-Stage1 $+$ Stage2 155.92
-Stage1 $+$ Stage2 $+$ Stage3 (SweetTokenizer) 44.35
-Table 4: Ablation study of spatial-temporal decoupling and progessive training schedule.
-UCF-101 rFVD $\downarrow$
-Baseline (w/o LLC) 67.94
-$+$ Spatial LLC 55.11
-$+$ Spatial & Temporal LLC (SweetTokenizer) 44.35
-$+$ CLIP [31]-based LLC 45.56
-$+$ Qwen [1]-based LLC (SweetTokenizer) 44.35
-ImageNet rFID $\downarrow$
-Baseline (w/o LLC) 1.26
-$+$ Spatial LLC (SweetTokenizer) 0.73
-Table 5: Ablation study of language-based latent codebook (LLC) settings on UCF-101 and ImageNet.
-Language-based Latent Codebook.
-We evaluate the impact of our architecture design on model performance, focusing on the effects of the spatial-temporal language-based latent codebooks (LLC) on the ImageNet and UCF-101 datasets.
-As illustrated in Table 5, the integration of the spatial language codebook consistently enhances performance across both video and image modalities, improving the rFVD score from 67.94 to 55.11 and the rFID score from 1.26 to 0.73.
-Moreover, our temporal language codebook significantly benefits video reconstruction tasks, further reducing the rFVD score from 55.11 to 44.35.
+Decoupled Query AutoEncoder.
+We demonstrate the effectiveness of our spatial-temporal decoupling design for token compression.
+As shown in Table 4, naively downsampling video tokens by 1D linear-interpolation from 5,120 patches into 1,280 tokens results in poor performance of a rFVD of 227.65.
+Training SweetTok without decoupling (similary to vanilla query-based method in LARP [39]) results in sub-optimal result, obtaining a rFVD of 35.15.
+Our proposed decoupling query strategy (DQAE) achieves best result of 20.46 rFVD.
+We attribute this improvement to two factors: (1) the flattening operation discards substantial consecutive temporal information, and (2) without decoupling, the model struggles to learn efficiently from the intertwined temporal and spatial information.
+Motion-enhanced Language Codebook.
+We evaluate the impact of our architecture design on model performance, focusing on the effects of the motion-enhanced language codebooks (MLC) on UCF-101 datasets.
+As illustrated in Table 5, vanilla language codebook design enhances rFVD performance from 29.45 to 24.80, leading to a sub-optimal result.
+Notably, our motion-enhanced temporal language codebook significantly benefits video reconstruction tasks, further reducing the rFVD score from 24.80 to 20.46.
 This underscores the importance of our unique design for video modalities.
-We attribute these improvements to the robust semantic features provided by language-based embeddings and their inherent alignment between visual and text spaces.
-Additionally, we compare different types of language-based embeddings, such as using CLIP [31] embeddings in place of Qwen-2.5B [1] embeddings.
-The experiments indicate that while CLIP embeddings yield satisfactory results, employing a more powerful pre-trained language model leads to superior performance.
+Additionally, we compare different types of language-based embeddings, such as using Qwen-2.5B [1] embeddings in place of CLIP [27] embeddings.
+The experiments indicate that naively increasing the complexity of pre-trained language model is not cost-effective for SweetTok.
 4.5 Visual Semantic Comprehension
 Few-Shot Visual Classification.
-To evaluate the semantic capabilities of SweetTokenizer, we conducted experiments on few-shot image classification and video action recognition tasks.
-In both experiments, we initially extracted visual tokens using SweetTokenizer and transformed them into natural language words via our LLM codebook.
+To evaluate the semantic capabilities of SweetTok, we conducted experiments on few-shot image classification and video action recognition tasks.
+In both experiments, we initially extracted visual tokens using SweetTok and transformed them into natural language words via our language-based codebook.
 Subsequently, we employed CLIP to compute the similarity between the visual inputs and text embeddings.
 The top 21 tokens with the highest similarity were selected to form a prompt for prediction using the Qwen LLM.
-For the image classification task, we adhered to the V2L protocol [59], comparing SweetTokenizer against other language-based visual tokenizers SPAE [54] and V2L.
+For the image classification task, we adhered to the V2L protocol [54], comparing SweetTok against other language-based visual tokenizers SPAE [49] and V2L.
 In the video action recognition task, we used ARN [53] and HF-AR [20] as baselines.
-The results in Table 6 indicate that SweetTokenizer achieved an accuracy of 90.8% on the miniImageNet dataset, surpassing SPAE 89.9% and V2L 87.6%.
-On the UCF-101 dataset, SweetTokenizer attain an average accuracy of 90.1%, outperforming ARN 83.1% and HF-AR 86.4%.
-These findings demonstrate SweetTokenizer robust semantic understanding and superior performance in both image and video tasks.
+The results in Table 6 indicate that SweetTok achieved an accuracy of 90.8% on the miniImageNet dataset, surpassing SPAE 89.9% and V2L 87.6%.
+On the UCF-101 dataset, SweetTok attain an average accuracy of 90.1%, outperforming ARN 83.1% and HF-AR 86.4%.
+These findings demonstrate SweetTok robust semantic understanding and superior performance in both image and video tasks.
+We also visualize the effectiveness of our MLC spatial-temporal words for video understanding in the supplementary.
+Methods rFVD $\downarrow$
+Baseline (w/o LC) 29.45
+$+$ LC 24.80
+$+$ MLC (SweetTok) 20.46
+$+$ CLIP [27]-based MLC (SweetTok) 20.46
+$+$ Qwen [1]-based MLC 20.12
+Table 5: Ablation study of different codebooks, including vanilla language codebook (LC), motion-enhanced language codebook (MLC) and more advanced pre-trained language codebook.
 Methods ImageNet UCF-101
@@
 HF-AR [20] - - - - 86.4
-SweetTokenizer 86.8 90.5 95.2 90.8 90.1
+SweetTok 86.8 90.5 95.2 90.8 90.1
 Table 6: Few-shot visual classification accuracy ($\uparrow$), evaluated on both image and video modality.
 5 Conclusions
-We present SweetTokenizer, an efficient visual tokenization framework that compresses spatial and temporal information through the cross-attention query autoencoder.
-Combined with language-based latent spaces and progressive training schedule, SweetTokenizer reduces token count by 4$\times$ for video data while maintaining high reconstruction fidelity.
-Our approach offers a compact representation of visual data, making it well-suited for downstream tasks such as visual generation, visual recognition, marking a significant step in efficient visual tokenization.
+We present SweetTok, an efficient video tokenization framework that compresses spatial and temporal information through the decoupled query autoencoder.
+Combined with motion-enhanced language codebook, SweetTok reduces token count for video data more effectively, achieving higher reconstruction fidelity compared to previous state-of-the-arts.
+Our approach offers a compact representation of video data, making it well-suited for downstream tasks such as video generation, understanding, marking a significant step in efficient video tokenization.
 Supplementary Material
 Notations Explanations
-$CQAE_{s,t}$ Cross-attention query autoencoder
+$DQAE_{s,t}$ Decoupled query autoencoder
+$\mathcal{P}$ Patchify Module
 $\mathcal{E}$ Encoder
@@
 $p_{t,h,w}$ Downsample ratio
-$\Omega_{s,t,m}$ Continuous latent query tokens
-$z_{s,t,m}$ Continuous latent query token
-$\hat{\Omega}_{s,t}$ Quantized latent query tokens
-$\hat{z}_{s,t}$ Quantized latent query token
+$\mathbf{Q}_{s,t,m}$ Continuous latent query tokens
+$\mathbf{Z}_{\mathbf{Q}_{s,t,m}}$ Continuous latent query token
+$\tilde{\mathbf{Q}}_{s,t}$ Quantized latent query tokens
+$\tilde{\mathbf{Z}}_{\mathbf{Q}_{s,t,m}}$ Quantized latent query token
 $v_{s,t}$ Continuous visual feature
-$\hat{v}_{s,t}$ Discrete visual feature
-$h(\cdot)$ Hidden unit
+$\tilde{v}_{s,t}$ Discrete visual feature
 $\Delta v$ Difference of consecutive features
@@
 $c_{adj,noun,adverb,verb}$ Codebook text embedding
-$\mathcal{P}$ Projector network
+$\mathcal{F}$ Projector network
 $sg(\cdot)$ Stop gradient operation
-$\Psi_{e,d}(\cdot)$ Pretrained encoder & decoder
-$\mathcal{L}_{c}$ Cross-Entropy loss
-$\mathcal{L}_{vq}$ Commitment loss
-$\mathcal{L}_{m}$ Mean-squared loss
-$\mathcal{L}_{g}$ Gan loss
-$\mathcal{L}_{p}$ Perceptual loss
 Table 7: Explanations for the notations in the main paper.
@@
 Then the video clip will be patchified to a resolution of 1 $\times$ 32 $\times$ 32 spatial feature and 4 $\times$ 32 $\times$ 32 temporal feature as illustrated in the main paper.
-The encoder $\mathcal{E}$ and decoder $\mathcal{D}$ in our tokenizer are both composed of 8 $CQAE_{s}$ modules and 4 $CQAE_{t}$ modules with 512 hidden states and 8 attention heads.
+The encoder $\mathcal{E}$ and decoder $\mathcal{D}$ in our tokenizer are both composed of 8 $DQAE_{s}$ modules and 4 $DQAE_{t}$ modules with 512 hidden states and 8 attention heads.
 Each modules consists of self-attention, feed-forward and cross-attention layers.
-Before the attention computation, the visual features will be reshaped into $[(BT)\times(HW)\times D]$ and $[(BHW)\times T\times D]$ for $CQAE_{s}$ and $CQAE_{t}$ modules, respectively.
+Before the attention computation, the visual features will be reshaped into $[(BT)\times(HW)\times D]$ and $[(BHW)\times T\times D]$ for $DQAE_{s}$ and $DQAE_{t}$ modules, respectively.
 The encoder $\mathcal{E}$ generates 256 spatial and 1024 temporal continuous latent tokens.
 These tokens are then passed to the quantizer $\mathcal{Q}$, which produces the quantized spatial and temporal latent tokens.
-The quantizer $\mathcal{Q}$ is composed of a spatial and temporal codebook and a GCN with two hidden layers with hidden dimension of 512 as the projector network $\mathcal{P}$.
+The quantizer $\mathcal{Q}$ is composed of a spatial and temporal codebook and a GCN with two hidden layers with hidden dimension of 512 as the projector network $\mathcal{F}$.
 To improve the training stability of the visual tokenizer, we adopt exponential moving average (EMA) updates with weight of 0.999 following [51].
@@
 We utilize VideoGPT following [41] as the default large language model for the video generative pre-training. All settings follow the protocol of [41].
-Tokenizer #Params #Tokens FID $\downarrow$
-VQVAE-2 [32] 13.5B 1024 31.11
-VQGAN [11] 1.4B 1024 74.3
-RQ-Transformer [26] 821M 1024 13.11
-ViT-VQGAN [46] 650M 1024 8.81
-OmniTokenizer [45] 650M 1024 7.45
-SweetTokenizer 650M 256 5.48
-Table 8: Comparison of class-conditional generation results on ImageNet-1K.
-Each image is of resolution of 256 $\times$ 256.
-We only compare autoregressive generators (AR) as baselines.
-“#Params” refers the parameter number of generator.
+Tokenizer PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+UCF-101
+LARP [39] 27.88 - 0.085
+SweetTok 29.27 0.7766 0.070
+ImageNet
+LlamaGen-16 [31] 20.79 0.675 -
+TokenFlow [25] 21.41 0.687 -
+LlamaGen-8 [31] 24.45 0.813 -
+SweetTok 30.23 0.826 0.068
+Table 8: More evaluation results on UCF-101 and ImageNet.
 Configuration Language Model Tokenizer
-Stage 1 Stage 2 Stage 3
-LLM init VideoGPT - - -
-Optimizer AdamW AdamW AdamW AdamW
+Image Finetune Video Training
+LLM init VideoGPT - -
+Optimizer AdamW AdamW AdamW
 Optimizer Hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.96$ $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1e^{-8}$ $\beta_{1}=0.9$, $\beta_{2}=0.999$
-Batch size per GPU 4 32 8 12
-Peak learning rate $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
-Discriminator peak learning rate - - - $1e^{-4}$
-Learning rate schedule Cosine Cosine Cosine Cosine
-Training steps 1000K 1000K 500K 500K
-Discriminator start steps - - - 20K
-Warm-up steps 10K 10K 10K 10K
-Weight decay 0.03 $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
-Numerical precision float16 float16 float16 bfloat16
-Table 9: The detailed training hyperparameters of SweetTokenizer.
+Batch size per GPU 4 32 12
+Peak learning rate $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
+Discriminator peak learning rate - - $1e^{-4}$
+Learning rate schedule Cosine Cosine Cosine
+Training steps 1000K 500K 1000K
+Discriminator start steps - - 20K
+Warm-up steps 10K 10K 10K
+Weight decay 0.03 $1e^{-4}$ $1e^{-4}$
+Numerical precision float16 float16 bfloat16
+Table 9: The detailed training hyperparameters of SweetTok.
 6.2 Training Datasets
@@
 6.4 Training Settings
-The detailed training hyper-parameter settings for SweetTokenizer are reported in Table 9.
+The detailed training hyper-parameter settings for SweetTok are reported in Table 9.
+During video training, we train the first 500K steps using proxy code following [51] to accelerate training.
+The remaining 500K is trained without proxy code.
 7 Additional Results
-7.1 Generation results for ImageNet
-Following the OmniTokenizer protocol [45], we train a VideoGPT-based generator on ImageNet-1K.
-As shown in Table 8, SweetTokenizer consistently outperforms all baselines in terms of FID.
-With fewer tokens, SweetTokenizer achieves 5.48 FID, outperforming OmniTokenizer by 26.4 %.
-These results align with findings on UCF-101, reinforcing that fewer tokens lead to better generation performance.
+7.1 More evaluation metrics
+We assess SweetTok using additional metrics: PSNR, SSIM, and LPIPS.
+As shown in Table 8, SweetTok outperforms all baselines on both video and image datasets, further validating the superiority of our model design.
 7.2 More Visualizations
@@
 The pixel-level differences between ground truth and model are shown, with brighter areas indicating greater disparity and darker areas reflecting consistency.
-As shown, SweetTokenizer exhibits fewer reconstruction differences compared to OmniTokenizer, demonstrating its superior performance.
-Fig 8 and Fig 9 visualize the reconstruction and generation results of SweetTokenizer on ImageNet-1K.
+As shown, SweetTok exhibits fewer reconstruction differences compared to OmniTokenizer, demonstrating its superior performance.
+Fig 9 visualize the reconstruction results of SweetTok on ImageNet-1K.
 For reconstruction, differences between models are highlighted in red blocks, with details shown in green blocks.
-Clearly, SweetTokenizer outperforms all baselines by a significant margin.
-Finally, we visualize the words from our LLM codebook in Fig 10, based on few-shot video action recognition tasks on the UCF-101 dataset.
+Clearly, SweetTok outperforms all baselines by a significant margin.
+Finally, we visualize the words from our MLC in Fig 10, based on few-shot video action recognition tasks on the UCF-101 dataset.
 We use adjectives, nouns, adverbs, and verbs as prompts to Qwen LLM for action prediction.
 Green and orange indicate meaningful words, while red marks meaningless ones.
-The visualization shows that correct verb words consistently lead to accurate predictions, even when other words are irrelevant, highlighting the importance of our adverb and verb LLM codebook modules for video action recognition.
-Figure 6: Comparison of the reconstruction results of OmniTokenizer and SweetTokenizer on UCF-101 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
-Figure 7: Comparison of the reconstruction results of OmniTokenizer and SweetTokenizer on K-600 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
-Figure 8: Comparison of the reconstruction results of TiTok, OmniTokenizer, and SweetTokenizer on ImageNet-1K dataset. Differences are selected by the red blocks and highlighted in the grean blocks.
-Figure 9: Class-conditional image generation results using SweetTokenizer, on ImageNet-1k dataset.
+The visualization shows that correct verb words consistently lead to accurate predictions, even when other words are irrelevant, highlighting the importance of our MLC modules for video action recognition.
+Figure 7: Comparison of the reconstruction results of OmniTokenizer and SweetTok on UCF-101 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
+Figure 8: Comparison of the reconstruction results of OmniTokenizer and SweetTok on K-600 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
+Figure 9: Comparison of the reconstruction results of TiTok, OmniTokenizer, and SweetTok on ImageNet-1K dataset. Differences are selected by the red blocks and highlighted in the grean blocks.
 Figure 10: Semantic words visualization for UCF-101. The visualization is based on few shot video action recognition tasks.
 8 Limitations
-Our tokenizer is not suitable for tasks requiring precise semantic understanding, like VQA, because the LLM codebook is trained in an unsupervised manner.
+Our tokenizer is not suitable for tasks requiring precise semantic understanding, like VQA, because the MLC is trained in an unsupervised manner.
 Without additional constraints, such as contrastive learning between image features from Qwen-VLM and text embeddings in our codebook, aligning the image and text domains is challenging.
-A promising direction for future work is to enhance SweetTokenizer into a semantically strong tokenizer by contrastive learning.
+A promising direction for future work is to enhance SweetTok into a semantically strong tokenizer by contrastive learning.